### v2

# Constrained mean-field control with singular control: Existence, stochastic maximum principle and constrained FBSDE

Abstract
This paper studies a class of mean-field control (MFC) problems with singular control under general dynamic state-control-law constraints. We first propose a customized relaxed control formulation to cope with the dynamic mixed constraints and establish the existence of an optimal control using compactification argument in the proper canonical spaces to accommodate the singular control. To characterize the optimal pair of regular and singular controls, we treat the controlled McKean-Vlasov process as an infinite-dimensional equality constraint and recast the MFC problem as an optimization problem on canonical spaces with constraints on Banach space, allowing us to derive the stochastic maximum principle (SMP) and a constrained BSDE using a novel Lagrange multipliers method. Additionally, we investigate the uniqueness and the stability result of the solution to the constrained FBSDE associated to the constrained MFC problem with singular control.
Keywords: Mean-field control, singular control, dynamic state-control-law constraint, compactification, stochastic maximum principle, Lagrange multiplier, constrained FBSDE
2020 MSC: 60H30, 49N80, 46N10, 60H10
1 Introduction
To overcome the high complexity in large stochastic system caused by the mutual influence among agents, mean-field game (MFG) problems (see [22] and [26]) has been proposed and popularized in academic studies, which is to find a Nash equilibrium in a competitive game formulation where each agent’s impact on the aggregate distribution is negligible. MFC problems has also gained considerable attention, which concerns the cooperative population where all agents share the same goal to attain the social optimum. The present paper targets at the MFC problem, which is often interpreted as an optimal control of McKean-Vlasov dynamics by a social planner. Some pioneer studies on this topic, just to name a few, can be found in [2, 7, 10, 1, 5] for the SMP method; [25] and [15] for the limit theory; [27] and [30] for the DPP and HJB equation on the Wasserstein space of probability measures; [4], [31] and [32] for the viscosity solution and comparison principle of the HJB equation. However, it is notable that most existing studies on MFC are limited to regular controls.
Singular control, also called bounded variation stochastic control, allows the displacement of the states to be discontinuous. In single agent’s models, optimal singular control problems have been extensively investigated. For instance, we refer to [29] for discussions on smooth fit principle; [20] for the existence of optimal singular controls based on the relaxed control formulation; [21] for the proof of DPP to account for singular controls; [3] for the SMP tailor-made for singular controls; [19] for a general theory on smooth fit principle. In the competitive game framework, there are a handful of studies on singular controls, see [11] for the n-player game with singular control as well as [17, 8, 9] for some MFG problems with singular control. However, theoretical studies for MFC problems with singular control are underexplored in the literature, and only a recent paper [14] studies the derivation of DPP and the quasi-variational inequality using two-layer parameterisations of stochastic processes.
The goal of the paper is to contribute to the study of MFC problems by featuring both the singular control and the joint dynamic state-control-law constraints. Motivated by practical risk regulations, liquidity requirements, or feasibility restraints, different types of constraints have been considered in stochastic control problems, but predominantly in single agent’s model. Recently, in mean-field models with only regular controls, there are some emerging research to incorporate state and/or law constraints, see [18] for the level-set approach to tackle the law constraint in MFC problems; [12] for the analysis of mean-field limits under state constraints; [13] for the formulation of controlled FP equation under state constraints; [6] for the optimization formulation of the constrained MFC problem and the derivation of the SMP using the generalized Lagrange multipliers method.
It is evident that the allowance of singular control in the mean-field model will entangle with the dynamic mixed constraints, which renders some existing methods for constrained MFC with regular controls not applicable in our framework. For example, it is difficult to generalize the level set approach in [18] to account for both control constraints and the singular control. Due to the possible discontinuity of the state process and its law under the singular control, characterizing the associated FP equation is still an open problem in the literature, and hence the approach in [13] is also not suitable to handle the singular control. On the other hand, some conventional methods that have been employed in unconstrained MFG and MFC problems with singular control such as DPP and compactification arguments also need to be carefully and substantially modified in the setting when pathwise constraints are imposed jointly on the state process and its distribution flow.
The present paper fills the gap in the literature by advancing some theoretical foundations for constrained MFC problems with singular control. The first contribution of the paper is to address the existence of an optimal pair of regular and singular controls by developing the customized relaxed control formulation and martingale problem to embed the dynamic mixed constraints. That is, we embed the dynamic mixed constraints in item (iv) of the admissible control rules in Definition 3.1. We then present in Lemma 3.5 a martingale problem characterization associated to the admissible control rules and establish in Lemma 3.7 the closedness property of the set of probability measures fulfilling the constraints. By developing technical compactification arguments to accommodate the singular control and combining the previous lemmas using the embedded constraints, we are able to show that the set of approximately optimal control rules is compact (see Lemma 3.8), and hence the set of optimal control rules is compact and convex that yields the existence of an optimal relaxed control (see Theorem 3.9). By standard argument, we also obtain the existence of an optimal strict control for the constrained MFC problem with singular control.
Our second contribution is to further characterize the optimal controls by establishing the SMP for the constrained MFC problem with singular control and derive the adjoint constrained BSDE. However, some conventional techniques such as the first order variation or the spike variation of controls may fail in our setting as the mixed constraints might be violated. Inspired by our recent study [6] for constrained MFC with regular controls, we propose to regard the controlled McKean-Vlasov dynamics with singular control as an infinite-dimensional equality constraint and recast the MFC problem as an optimization problem on some proper canonical spaces with general constraints on Banach space (see the formulation (4)-(17)). Similar to [6], we are allowed to derive the SMP by using the generalized Fritz-John (FJ) optimality condition and the Lagrange multipliers method. To be more precise, when the forward McKean-Vlasov SDE is taken as an equality constraint, the FJ condition for the control variable gives the SMP, and the Lagrange multipliers in the space of Itô processes induced by the FJ condition on the state variable yields the adjoint processes as the solution to the constrained BSDE, see Theorem 4.2.
As the third contribution of the present paper, we also analyze the non-standard constrained FBSDE in (5)-(27) associated to our constrained MFC problem with singular control. It is worth noting that the Lagrange multiplier $\eta$ in Theorem 4.2 is treated as part of the solution to the constrained BSDE problem. We then verify the constraints qualification and obtain the KKT optimality condition in Lemma 5.1. As a result, we further establish the uniqueness of the solution to the nonstandard constrained FBSDE in (5)-(27), see Theorem 5.4. Furthermore, under some enhanced assumptions, we also prove the stability result of the solution in the sense of (5.6) in Theorem 5.6.
The rest of the paper is organized as follows. Section 2 introduces the constrained MFC problem with singular control and standing assumptions imposed on the model. Section 3 gives the relaxed control formulation and establishes the existence of optimal control to the constrained MFC problem by applying compactification arguments. Section 4 formulates the equivalent optimization problem with constraints in Banach space, for which we develop the SMP and ajoint constrained BSDE using Lagrange multipliers method. Lastly, the uniqueness and the stability result of the solution to the constrained FBSDE associated to the constrained MFC problem with singular control are discussed in Section 5. The proofs of some auxiliary results are provided in Appendix.
Notations. We list below some notations that will be frequently used throughout the paper:
$(E^{*},\lVert\cdot\rVert_{E^{*}})$ Dual space of a Banach space $(E,\lVert\cdot\rVert_{E})$
$\langle\cdot,\cdot\rangle_{E,E^{*}}$ Pairing between Banach space $(E,\lVert\cdot\rVert_{E})$ and its dual
$a\geq_{K}0~(a>_{K}0)$ $a\in K~(a\in K^{\mathrm{o}})$, with $K$ a closed (linear) cone
DfDf Fréchet derivative of $f:E\to\mathbb{R}$
$\nabla_{x}f$ Gradient of $f:\mathbb{R}^{n}\to\mathbb{R}$
$D_{X}f$ Partial Fréchet derivative of $f$ with respect to $X$
$D_{\alpha}f$ Partial Fréchet derivative of $f$ with respect to $\alpha$
$L^{p}((A,\mathscr{B}(A),\lambda_{A});E)$ Set of $L^{p}$-integrable $E$-valued mappings on measure space
$(A,\mathscr{B}(A),\lambda_{A})$; write $L^{p}(A;E)$ for short
$L^{p}((A,\mathscr{B}(A),\lambda_{A});E)/\mathord{\sim}$ Quotient set of $L^{p}(A;E)$ under relation $f\mathord{\sim}g$ iff $f=g$ $\lambda_{A}$-a.e.;
write $L^{p}(A;E)/\mathord{\sim}$ for short
$\mathbb{E}$ ($\mathbb{E}^{\prime}$) Expectation operator under probability measure $\mathbb{P}$ ($\mathbb{P}^{\prime}$)
${\cal P}_{p}(E)$ Set of probability measures on $E$ with finite $p$-order moments
$\mathcal{M}[0,T]$ ($\mathcal{M}^{+}[0,T]$) Set of signed (non-negative) Radon measures on $[0,T]$
$\mathrm{BV}^{k}[0,T]$ Set of functions $f:[0,T]\to\mathbb{R}^{k}$ with bounded variation
$D^{k}[0,T]$ Set of càdlàg functions $f:[0,T]\to\mathbb{R}^{k}$
2 Problem Formulation
2.1 MFC with singular control and mixed constraints
Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered probability space with the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ satisfying the usual conditions. For $r,l,k\in\mathbb{N}$ and $p>2$, let $W=(W_{t})_{t\in[0,T]}$ be a standard $r$-dimensional $(\mathbb{P},\mathbb{F})$-Brownian motion and $U\subset\mathbb{R}^{l}$ be a closed and convex (regular) set as the policy space. Define the space of regular controls $\mathscr{U}[0,T]$ as the set of $\mathbb{F}$-progressively measurable processes such that $\mathbb{E}[\int_{0}^{T}|\alpha_{t}|^{p}\mathrm{d}t]<\infty$. For the space of singular controls, define $A^{k}[0,T]$ as the set of functions $a:[0,T]\mapsto\mathbb{R}^{k}$ such that $a=(a_{t}^{i})_{t\in[0,T]}$ is càdlàg and non-decreasing with $a_{0}^{i}=0$ for $i=1,\dots,k$. Let $\mathcal{A}^{k}[0,T]$ be the set of $\mathbb{F}$-progressively measurable processes $\zeta=(\zeta_{t})_{t\in[0,T]}$ with paths in $A^{k}[0,T]$ and $\mathbb{E}[|\zeta_{T}|^{p}]<\infty$.
Let $(b,\sigma):[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\mapsto\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ be measurable mappings and $G:[0,T]\mapsto\mathbb{R}^{n\times k}$ be a continuous function. For $(\alpha,\zeta)=(\alpha_{t},\zeta_{t})_{t\in[0,T]}\in\mathscr{U}[0,T]\times\mathcal{A}^{k}[0,T]$, we consider the controlled McKean-Vlasov SDE that, for $t\in(0,T]$,
$\mathrm{d}X_{t}^{\alpha,\zeta}$ $=b(t,X_{t}^{\alpha,\zeta},\mathcal{L}(X_{t}^{\alpha,\zeta}),\alpha_{t})\mathrm{d}t+\sigma(t,X_{t}^{\alpha,\zeta},\mathcal{L}(X_{t}^{\alpha,\zeta}),\alpha_{t})\mathrm{d}W_{t}+G(t)\mathrm{d}\zeta_{t},$ (1)
and $X_{0}^{\alpha,\zeta}=\xi\in L^{p}((\Omega,\mathcal{F}_{0},\mathbb{P});\mathbb{R}^{n})$ with the law of $\xi$ under $\mathbb{P}$ being $\mathcal{L}(\xi)=\nu\in\mathcal{P}_{p}(\mathbb{R}^{n})$.
For the above controlled McKean-Vlasov dynamics, we introduce dynamic state-control-law constraints. Let $d\in\mathbb{N}$ be the number of inequality constraints. For $i\in I:=\{1,\dots,d\}$, consider $\phi^{i}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\mapsto\mathbb{R}$ with $i\in I$ such that the dynamic state-control-law inequality constraints are described by, $\mathrm{d}t\otimes\mathrm{d}\mathbb{P}$-a.s.
$\phi^{i}(t,X_{t}^{\alpha,\zeta},\mathcal{L}(X_{t}^{\alpha,\zeta}),\alpha_{t})\geq 0,$ (2)
where $X^{\alpha,\zeta}=(X_{t}^{\alpha,\zeta})_{t\in[0,T]}$ obeys the McKean-Vlasov SDE (1). Denote by $\mathscr{U}_{\rm ad}[0,T]$ the admissible control set of processes $(\alpha,\zeta)\in\mathscr{U}[0,T]\times\mathcal{A}^{k}[0,T]$ such that there exists a càdlàg $\mathbb{F}$-adapted ($\mathbb{R}^{n}$-valued) process $X^{\alpha,\zeta}=(X_{t}^{\alpha,\zeta})_{t\in[0,T]}$ satisfying (1) and (2).
The goal of the MFC problem is to minimize the following cost functional that, for $(\alpha,\zeta)=(\alpha_{t},\zeta_{t})_{t\in[0,T]}\in\mathscr{U}_{\rm ad}[0,T]\neq\varnothing$,
$\inf_{(\alpha,\zeta)\in\mathscr{U}_{\rm ad}[0,T]}J(\alpha,\zeta):=\inf_{(\alpha,\zeta)\in\mathscr{U}_{\rm ad}[0,T]}\mathbb{E}\left[\int_{0}^{T}f(t,X_{t}^{\alpha,\zeta},\mathcal{L}(X_{t}^{\alpha,\zeta}),\alpha_{t})\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}\right],$ (3)
where $f:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\mapsto\mathbb{R}$ and $c:[0,T]\mapsto\mathbb{R}^{k}$ are respectively the running cost function and the cost function associated with singular controls.
2.2 Standing assumptions
Let us first unpack some standing assumptions imposed on the model coefficients, which will be used in different sections.
Assumption 1.
  • (A1)
    The mappings $b,\sigma,f$ and $\phi^{i}$, for $i\in I$ are Borel measurable, $b,\sigma$ are continuous and $f$ is lower semi-continuous (l.s.c.) in $(x,\mu,u)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$.
  • (A2)
    For $i\in I$, $\phi^{i}$ is upper semi-continuous (u.s.c.) in $(x,\mu,u)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$, and there exists $M>0$ such that $|\phi^{i}(t,x,\mu,u)|\leq M(1+|x|+\sqrt{M_{2}(\mu)}+|u|)$ for $(i,x,u)\in I\times\mathbb{R}^{n}\times\mathbb{R}^{l}$ and $M_{2}(\mu):=\int_{\mathbb{R}^{n}}|x|^{2}\mu(\mathrm{d}x)$. Moreover, $\phi^{i}$ is uniformly u.s.c. in $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$ w.r.t. $(t,x,u)\in[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}$.
  • (A3)
    The mappings $b(t,x,\mu,u)$ and $\sigma(t,x,\mu,u)$ are uniformly Lipschitz continuous in $(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$, i.e., there is $M>0$ independent of $u\in\mathbb{R}^{l}$ such that, for all $(x,\mu),(x^{\prime},\mu^{\prime})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$ and $(t,u)\in[0,T]\times\mathbb{R}^{l}$,
    $\left|b(t,x^{\prime},\mu^{\prime},u)-b(t,x,\mu,u)\right|+\left|\sigma(t,x^{\prime},\mu^{\prime},u)-\sigma(t,x,\mu,u)\right|\leq M(|x-x^{\prime}|+\mathcal{W}_{2,\mathbb{R}^{n}}(\mu,\mu^{\prime})).$
  • (A4)
    There is $M>0$ such that $|b(t,x,\mu,u)|+|\sigma(t,x,\mu,u)|\leq M(1+|x|+\sqrt{M_{2}(\mu)}+|u|)$ for $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$.
  • (A5)
    $G:[0,T]\mapsto\mathbb{R}^{n\times k}$ and $c:[0,T]\mapsto[0,\infty)^{k}$ are deterministic continuous functions. Moreover, there is $c_{0}>0$ such that $c^{i}(t)\geq c_{0}$ for all $t\in[0,T]$.
  • (A6)
    There exist $M,M_{1}>0$ such that, for all $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$,
    $M(1+|x|^{p}+\sqrt{M_{2}(\mu)})+M_{1}|u|^{p}\leq f(t,x,\mu,u)\leq M(1+|x|^{p}+\sqrt{M_{2}(\mu)}+|u|^{p}).$
Assumption 2.
  • (B1)
    The mappings $b,\sigma,f$ and $\phi^{i}$ are Borel measurable and continuously differentiable w.r.t. $x\in\mathbb{R}^{n}$. Moreover, $\phi^{i}$ is jointly continuous in $(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$.
  • (B2)
    There exists $M>0$ such that, for all $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$,
    $\sum_{i\in I}\left(|\nabla_{x}\phi^{i}(t,x,\mu,u)|+|\nabla_{u}\phi^{i}(t,x,\mu,u)|\right)$ $\leq M.$
  • (B3)
    The condition (A3) in Assumption 1 holds.
  • (B4)
    There exists $M>0$ such that $|\delta(t,x,\mu,u)|\leq M(1+|x|+|u|+\sqrt{M_{2}(\mu)})$ for $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$ and $\delta\in\{b,\sigma,\nabla_{u}b(\cdot),\nabla_{u}\sigma(\cdot),\nabla_{x}f(\cdot),\nabla_{u}f(\cdot)\}$.
  • (B5)
    The mappings $b(t,x,\mu,u),\sigma(t,x,\mu,u),f(t,x,\mu,u)$ and $\phi^{i}(t,x,\mu,u)$ are continuously Lions-differentiable w.r.t. $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$. Moreover, there exists $M>0$ such that, for all $(t,x,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$,
    $\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\mu}f(t,x,\mu,u)(x^{\prime})|^{2}\mu(\mathrm{d}x^{\prime})\right)^{\frac{1}{2}}\leq M\left(1+|x|+|u|+\sqrt{M_{2}(\mu)}\right),$
    and moreover $\sum_{i\in I}\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\mu}\phi^{i}(t,x,\mu,u)(x^{\prime})|^{2}\mu(\mathrm{d}x^{\prime})\right)^{\frac{1}{2}}\leq M$.
  • (B6)
    $G:[0,T]\mapsto\mathbb{R}^{n\times k}$ and $c:[0,T]\mapsto[0,\infty)^{k}$ are deterministic continuous functions.
Assumption 3.
  • (C1)
    The constrained function $\phi$ is independent of the control variable and has the linear form $\phi(t,x,\mu)=A(t)x+B(t)\int_{\mathbb{R}^{n}}x\mu(\mathrm{d}x)+C(t)$ for $(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$. Here, $A=(A(t))_{t\in[0,T]}$, $B=(B(t))_{t\in[0,T]}$ and $C=(C(t))_{t\in[0,T]}\in C([0,T];\mathbb{R}^{n})$ and $|A(t)|>0$ and $|A(t)+B(t)|>0$ for $t\in[0,T]$.
  • (C2)
    The Hamiltonian $H$ defined by $H(t,x,\mu,u,y,z):=b(t,x,\mu,u)^{\top}y+\mathrm{tr}(\sigma(t,x,\mu,u)^{\top}z)+f(t,x,\mu,u)$ is strictly convex in $u\in U$. There exists a measurable mapping $\hat{u}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto U\subset\mathbb{R}^{l}$ with at most quadratic growth in $(x,\mu,y,z)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ such that $\hat{u}(t,x,\mu,y,z)\in\mathop{\arg\min}_{u\in U}H(t,x,\mu,u,y,z)$ and $\hat{u}$ is Lipschitz continuous, i.e., there is $M>0$ such that, for $(x_{i},\mu_{i},y_{i},z_{i})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ with $i=1,2$,
    $|\hat{u}(t,x_{1},\mu_{1},y_{1},z_{1})-\hat{u}(t,x_{2},\mu_{2},y_{2},z_{2})|$
    $\qquad\qquad\leq M(|x_{1}-x_{2}|+\mathcal{W}_{2,\mathbb{R}^{n}}(\mu_{1},\mu_{2})+|y_{1}-y_{2}|+|z_{1}-z_{2}|).$
    Set $\hat{b}(t,x,\mu,y,z):=b(t,x,\mu,\hat{u}(t,x,\mu,y,z))$, $\hat{\sigma}(t,x,\mu,y,z):=\sigma(t,x,\mu,\hat{u}(t,x,\mu,y,z))$ and $\hat{f}(t,x,\mu,y,z):=f(t,x,\mu,\hat{u}(t,x,\mu,y,z))$.
  • (C3)
    Define $F(t,x,\mu,u,y,z):=\nabla_{x}H(t,x,\mu,u,y,z)+\int_{\mathbb{R}^{n}}\partial_{\mu}H(t,x^{\prime},\mu,u,y,z)(x)\mu(\mathrm{d}x^{\prime})$,
    $\hat{F}(t,x,\mu,y,z):=F(t,x,\mu,\hat{u}(t,x,\mu,y,z),y,z)$ and $\Phi(t,x,\mu,y,z):=(-\hat{F},\hat{b},\hat{\sigma})(t,x,\mu,y,z)$. The following monotone condition holds that
    $(\Phi(t,x_{1},\mu,y_{1},z_{1})-\Phi(t,x_{2},\mu,y_{2},z_{2}))(x_{1}-x_{2},y_{1}-y_{2},z_{1}-z_{2})$
    $\qquad+\beta(|x_{1}-x_{2}|^{2}+|y_{1}-y_{2}|^{2}+|z_{1}-z_{2}|^{2})\leq 0,$
    where $\beta>\frac{M}{2}$ with $M$ being given in (A3). The inner product is read as
    $(\Phi(t,x_{1},\mu,y_{1},z_{1})-\Phi(t,x_{2},\mu,y_{2},z_{2}))(x_{1}-x_{2},y_{1}-y_{2},z_{1}-z_{2})$
    $\qquad=-(\hat{F}(t,x_{1},\mu,y_{1},z_{1})-\hat{F}(t,x_{2},\mu,y_{2},z_{2})^{\top}(x_{1}-x_{2})$
    $\qquad\quad+(\hat{b}(t,x_{1},\mu,y_{1},z_{1})-\hat{b}(t,x_{2},\mu,y_{2},z_{2}))^{\top}(y_{1}-y_{2})$
    $\qquad\quad+\mathrm{tr}\left((\hat{\sigma}(t,x_{1},\mu,y_{1},z_{1})-\hat{\sigma}(t,x_{2},\mu,y_{2},z_{2}))^{\top}(z_{1}-z_{2})\right).$
  • (C4)
    $\hat{b}$, $\hat{\sigma}$, $\hat{F}$ are Lipschitz continuous, i.e., there is $M>0$ such that, for $(x_{i},\mu_{i},y_{i},z_{i})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ with $i=1,2$, $\delta\in\{\hat{b},\hat{\sigma},\hat{F}\}$ and $x\in\mathbb{R}^{n}$,
    $|\delta(t,x_{1},\mu_{1},y_{1},z_{1})-\delta(t,x_{2},\mu_{2},y_{2},z_{2})|$
    $\qquad\leq M(|x_{1}-x_{2}|+\mathcal{W}_{2,\mathbb{R}^{n}}(\mu_{1},\mu_{2})+|y_{1}-y_{2}|+|z_{1}-z_{2}|).$
Assumption 2 strengthens Assumption 1 by requiring the differentiability of the coefficients and imposing growth conditions on their derivatives, while it is weaker than Assumption 1 by relaxing the growth condition on the running cost function $f$ and omitting the requirement for a uniform positive constant $c_{0}$ of the cost function $c$. Assumption 3 introduces the monotonicity to the coefficients and guarantees the existence of a Lipschitz continuous minimizer for the Hamiltonian.
3 Existence of Optimal Controls
3.1 Relaxed control formulation
To confirm the existence of an optimal pair of regular and singular controls to problem (3), we first reformulate the problem in a relaxed control framework such that the compactification method can be utilized. To this end, let us introduce the following basic spaces:
  • The space $D^{n}[0,T]$ is endowed with the Skorokhod metric $d_{D}$ and the corresponding Borel $\sigma$-algebra denoted by $\mathcal{F}^{D}$. Let $\mathcal{F}_{t}^{D}$ be the Borel $\sigma$-algebra up to time $t$ (c.f. [20]) for $t\in[0,T]$.
  • The space $\mathcal{Q}$ of relaxed controls is defined as the set of measures $q$ in $0,T]\times U$ with the first marginal equal to the Lebesgue measure and $\int_{[0,T]\times U}|u|^{p}q(\mathrm{d}t,\mathrm{d}u)<\infty$. We endow the space $\mathcal{Q}$ with the $2$-Wasserstein metric on $\mathcal{P}_{2}([0,T]\times U)$ given by $d_{\mathcal{Q}}(q^{1},q^{2})=\mathcal{W}_{2,[0,T]\times U}\left(\frac{q^{1}}{T},\frac{q^{2}}{T}\right)$, where the metric on $0,T]\times U$ is given by $((t_{1},u_{1}),(t_{2},u_{2}))\mapsto|t_{2}-t_{1}|+|u_{2}-u_{1}|$. Note that each $q\in\mathcal{Q}$ can be identified with a measurable function $0,T]\in t\mapsto q_{t}\in\mathcal{P}_{2}(U)$, defined uniquely up to $\mathrm{a.s.}$ by $q(\mathrm{d}t,\mathrm{d}u)=q_{t}(\mathrm{d}u)\mathrm{d}t$. In the sequel, we will always refer to the measurable mapping $q=(q_{t})_{t\in[0,T]}$ to a relaxed control in $\mathcal{Q}$. Let $\mathcal{F}^{\mathcal{Q}}$ be the Borel $\sigma$-algebra of $\mathcal{Q}$ and $\mathcal{F}_{t}^{\mathcal{Q}}$ be the $\sigma$-algebra generated by the maps $q\mapsto q([0,s]\times V)$ with $s\in[0,t]$ and Borel measurable $V\subset U$.
  • The space $A^{k}[0,T]$ is a subset of $\mathcal{V}\subset\mathrm{BV}^{k}[0,T]$, where the set $\mathcal{V}\subset\mathrm{BV}^{k}[0,T]$ is the collection of mappings $a:[0,T]\to\mathbb{R}^{k}$ such that $a$ is càdlàg and $a(0)=0$. As in [20], there is a natural one-to-one correspondence between $\mathcal{V}$ and $(\mathcal{M}[0,T])^{k}$, that is, for $a=(a^{1},\ldots,a^{k})\in\mathcal{V}$ and $i=1,\ldots,k$, define $\nu_{i}((s,t])=a^{i}(t)-a^{i}(s)$ for $0\leq s\leq t\leq T$. Then, $\mathcal{V}$ is the dual space of $(C[0,T])^{k}$, and hence we endow $\mathcal{V}$ with the corresponding weak topology (which is also equivalent to the Fortet-Mourier norm on ($\mathcal{M}[0,T])^{k}$, c.f. [5]). Thus, $A^{k}[0,T]$ is a closed subset of $(\mathcal{M}[0,T])^{k}$ and can also be viewed as $(\mathcal{M}^{+}[0,T])^{k}$. Let $\mathcal{F}^{A}$ be the Borel $\sigma$-algebra and $\mathcal{F}^{A}_{t}$ be the Borel $\sigma$-algebra defined up to time $t\in[0,T]$ (cf. [20]).
Now, we introduce the sample space considered in the paper, which is given by
$\Omega:=\widehat{\Omega}\times A^{k}[0,T]=D^{n}[0,T]\times\mathcal{Q}\times A^{k}[0,T].$ (4)
[…]
$d(R_{1},R_{2})=\mathcal{W}_{2,\widehat{\Omega}}(R_{1}|_{\widehat{\Omega}},R_{2}|_{\widehat{\Omega}})+d_{\rm FM,\Omega}(R_{1},R_{2}),$
[…]
Next, we give the definition of relaxed controls. In particular, one important task is to guarantee that the dynamic state-control-law constraints are fulfilled when applying the compactification argument. For this purpose, we propose to embed the constraints into the definition of admissible control rules in the following sense. Recall that $\nu\in\mathcal{P}_{p}(\mathbb{R}^{n})$ is the law of initial state in SDE (1).
In the sequel, we will always consider the weak formulation, $\mathrm{i.e.}$, we will take $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ introduced in this subsection to be the canonical space.
Definition 3.1.
[…]
  • (i)
    $R\circ X_{0}^{-1}=\nu$;
  • (ii)
    $\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}|u|^{p}q_{t}(\mathrm{d}u)\mathrm{d}t+|\zeta_{T}|^{p}\right]<\infty$;
  • (iii)
    for all $g\in C_{b}^{2}(\mathbb{R}^{n})$, $M^{R}g=(M_{t}^{R}g)_{t\in[0,T]}$ is an $(R,\mathbb{F})$-martingale, where
    $M_{t}^{R}g$ $:=g(X_{t})-\int_{0}^{t}\int_{U}\mathbb{G}g(s,X_{s},R^{X}(s),u)q_{s}(\mathrm{d}u)\mathrm{d}s-\int_{0}^{t}\nabla_{x}g(X_{s-})G(s)\mathrm{d}\zeta_{s}$
    $\quad-\sum_{0<s\leq t}\left(g(X_{s})-g(X_{s-})-\nabla_{x}g(X_{s-})\Delta X_{s}\right)$
    with $R^{X}(s):=R\circ X_{s}^{-1}$, $\Delta X_{s}:=X_{s}-X_{s-}$ and
    $\mathbb{G}g(t,x,\mu,u):=b(t,x,\mu,u)\nabla_{x}g(x)+\frac{1}{2}\mathrm{tr}\left(\sigma\sigma^{\top}(t,x,\mu,u)\nabla_{xx}^{2}g(x)\right).$
  • (iv)
    for $i\in I$, $m\otimes R(A_{i}^{R})=T$ with the admissible set $A_{i}^{R}$ defined by
    $A_{i}^{R}=\left\{(t,\omega)\in[0,T]\times\Omega;~\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\geq 0\right\},$ (5)
    where $m\otimes R$ denotes the product measure of Lebesugue measure $m$ and $R$.
3.2 Existence of optimal relaxed controls
For $\nu\in\mathcal{P}_{p}(\mathbb{R}^{n})$, let $\mathcal{R}_{\rm ad}(\nu)$ be the set of admissible control rules $R$ in Definition 3.1. For $R\in\mathcal{R}_{\rm ad}(\nu)$, we have the following martingagle measure characterization:
Lemma 3.2.
$R\in\mathcal{R}_{\rm ad}(\nu)$ iff there is a filtered probability space $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{F}^{\prime}=(\mathcal{F}^{\prime}_{t})_{t\in[0,T]},P^{\prime})$ supporting an $n$-dimensional $\mathbb{F}^{\prime}$-adapted process $X=(X_{t})_{t\in[0,T]}$, $r$-dimensional orthogonal $\mathbb{F}^{\prime}$-martingale measures $N=(N^{1},\ldots,N^{r})$ on $U\times[0,T]$, each with $\mathcal{F}^{\prime}_{t}$-progressively measurable intensity $q_{t}(\mathrm{d}u)\mathrm{d}t$ and a $k$-dimensional $\mathbb{F}^{\prime}$-progressively measurable nondecreasing process $\zeta=(\zeta_{t})_{t\in[0,T]}$ such that $R=P^{\prime}\circ(X,q,\zeta)^{-1}$, and the following conditions hold:
  • (i)
    $P^{\prime}\circ X_{0}^{-1}=\nu$;
  • (ii)
    $\mathbb{E}^{P^{\prime}}\left[\int_{0}^{T}\int_{U}|u|^{p}q_{t}(\mathrm{d}u)\mathrm{d}t+|\zeta_{T}|^{p}\right]<\infty$;
  • (iii)
    the state dynamics obeys that, $P^{\prime}$-a.s.
    $\mathrm{d}X_{t}=\int_{U}b(t,X_{t},\mathcal{L}^{P^{\prime}}(X_{t}),u)q_{t}(\mathrm{d}u)\mathrm{d}t+\int_{U}\sigma(t,X_{t},\mathcal{L}^{P^{\prime}}(X_{t}),u)N(\mathrm{d}u,\mathrm{d}t)+G(t)\mathrm{d}\zeta_{t},$
    where $\mathcal{L}^{P^{\prime}}(X_{t})$ denotes the law of $X_{t}$ under $P^{\prime}$.
  • (iv)
    the constraints hold that, for $m\otimes P^{\prime}$-$\mathrm{a.s.}$, $\int_{U}\phi^{i}(t,X_{t},\mathcal{L}^{P^{\prime}}(X_{t}),u)q_{t}(du)\geq 0$ for $i\in I$.
Moreover, there exists $C>0$ depending on $M,\nu,p,T$ such that $\mathbb{E}^{P^{\prime}}[\sup_{t\in[0,T]}|X_{t}|^{p}]\leq C\left\{1+\mathbb{E}\left[|\xi|^{p}\right]\right\}$ with $M$ stated in Assumption 1.
Proof.
The first assertion follows from [16], and the second assertion can be easily proved by basic moment estimations under Assumption 1. ∎
Let us define the cost functional on $\mathcal{R}_{\rm ad}(\nu)$ by
$\Gamma(R):=\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}f(t,X_{t},R^{X}(t),u)q_{t}(\mathrm{d}u)\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}\right].$ (6)
By using Assumption 1, $v^{*}:=\inf_{R\in\mathcal{R}_{\rm ad}(\nu)}\Gamma(R)\in\mathbb{R}$. Then, the optimal control rule set is given by
$\mathcal{R}_{\rm opt}(\nu):=\inf\{R\in\mathcal{R}_{\rm ad}(\nu);~\Gamma(R)=v^{*}\}.$ (7)
Remark 3.3.
For any strict admissible control $(\alpha,\zeta)\in\mathscr{U}_{\rm ad}[0,T]$, let $X^{\alpha,\zeta}$ be the state process obeyed by (1) and $R$ be the joint law of $(X^{\alpha,\zeta}_{\cdot},\delta_{\alpha_{\cdot}}(\mathrm{d}u)\mathrm{d}t,\zeta_{\cdot})$. Then, $R$ can be easily verified to be an admissible control and $J(\alpha,\zeta)=\Gamma(R)$ with $J(\alpha,\zeta)$ defined in (3). This yields that $v^{*}\leq\inf_{(\alpha,\zeta)\in\mathscr{U}_{\rm ad}[0,T]}J(\alpha,\zeta)$.
To arrive at the reverse inequality in Remark 3.3, for $(t,x,\nu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$, define a subset of $\mathbb{R}^{n}\times\mathbb{R}^{n\times n}\times\mathbb{R}\times\mathbb{R}^{m}$ by
$K(t,x,\nu)$ (8)
$\quad:=\left\{(b(t,x,\nu,u),\sigma\sigma^{\top}(t,x,\nu,u),z,w);~z\geq f(t,x,\nu,u),w_{i}\leq\phi^{i}(t,x,\nu,u),~i\in I,u\in U\right\}.$
We make the following assumption on the set $K(t,x,\nu)$:
Assumption 4.
For any $(t,x,\nu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$, the set $K(t,x,\nu)$ is convex.
Then, we have the following lemma whose proof is provided in Appendix.
Lemma 3.4.
Let Assumption 4 hold. For any admissible (relaxed) control rule $R\in\mathcal{R}_{\rm ad}(\nu)$, there exists an admissible control $(\alpha,\zeta)\in\mathscr{U}_{\rm ad}[0,T]$ such that $J(\alpha,\zeta)\leq\Gamma(R)$.
We next study the existence of optimal relaxed controls. Consider the McKean-Vlasov SDE (1) and define $Y_{t}:=X_{t}^{\alpha,\zeta}-\int_{0}^{t}G(s)\mathrm{d}\zeta_{s}$ for $t\in[0,T]$. It follows from the dynamics (1) that
$\mathrm{d}Y_{t}=b(t,X_{t}^{\alpha,\zeta},\mathcal{L}(X_{t}^{\alpha,\zeta}),\alpha_{t})\mathrm{d}t+\sigma(t,X_{t}^{\alpha,\zeta},\mathcal{L}(X_{t}^{\alpha,\zeta}),\alpha_{t})\mathrm{d}W_{t},~~Y_{0}=\xi.$
Motivated by this expression, we have the next result whose proof is omitted because the necessary part is obvious and the sufficiency can be shown by the classical martingale argument.
Lemma 3.5.
[…]
  • (i)
    $Y$ has continuous trajectories and $R(X_{\cdot}=Y_{\cdot}+\int_{0}^{\cdot}G(s)\mathrm{d}\zeta_{s})=1$; (ii) $R\circ Y_{0}^{-1}=\nu$;
  • (iii)
    $\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}|u|^{p}q_{t}(\mathrm{d}u)\mathrm{d}t+|\zeta_{T}|^{p}\right]<\infty$; (iv) for any $g\in C_{b}^{2}(\mathbb{R}^{n})$, $\tilde{M}_{t}^{R}g$ is a $(R,\mathbb{F})$-martingale, where $\tilde{M}_{t}^{R}g:=g(Y_{t})-\int_{0}^{t}\int_{U}\tilde{\mathbb{G}}g(s,X_{s},R^{X}(s),u)(Y_{s})q_{s}(\mathrm{d}u)\mathrm{d}s$ with
    $\tilde{\mathbb{G}}g(t,x,\mu,u)(y):=b(t,x,\mu,u)\nabla_{x}g(y)+\frac{1}{2}\mathrm{tr}(\sigma\sigma^{\top}(t,x,\mu,u)\nabla_{xx}^{2}g(y)).$
  • (v)
    for $i\in I$, we have $m\otimes R(A_{i}^{R})=T$.
Moreover, we have
Lemma 3.6.
The objective functional $R\mapsto\Gamma(R)$ defined by (6) is l.s.c. on $\mathcal{R}_{\rm ad}(\nu)$.
Proof.
Let us define the mapping $\Lambda(R,\omega):\mathcal{P}_{2}(\Omega;\widehat{\Omega})\times\Omega\mapsto\mathbb{R}$ by
$\Lambda(R,\omega)=\int_{0}^{T}\int_{U}f(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}(\omega).$
[…]
$\liminf_{\ell\to\infty}\Gamma(R_{\ell})=\liminf_{\ell\to\infty}\mathbb{E}^{R_{\ell}}[\Lambda(R_{\ell},\omega)]$
$\qquad\geq\liminf_{\ell\to\infty}\left\{\mathbb{E}^{R_{\ell}}[\Lambda(R_{\ell},\omega)]-\mathbb{E}^{R_{\ell}}[\Lambda(R,\omega)]\right\}+\liminf_{\ell\to\infty}\mathbb{E}^{R_{\ell}}[\Lambda(R,\omega)]\geq\mathbb{E}^{R}[\Lambda(R,\omega)].$
This yields the desired result that the objective functional $\Gamma$ is l.s.c. on $\mathcal{R}_{\rm ad}(\nu)$. ∎
The next result also holds, whose proof is reported in Appendix.
Lemma 3.7.
[…]
We next show that the $\epsilon$-optimal set is compact.
Lemma 3.8.
[…]
Proof.
[…]
$\tilde{R}_{\ell}(F\times C)=\int_{F}\mathbf{1}_{C}(y(\omega))R_{\ell}(\mathrm{d}\omega),~~\forall F\in\mathcal{F},~C\in\mathcal{F}^{\mathcal{C}}$ (9)
with $y(\omega)$ being defined by $y_{\cdot}(\omega):=\psi_{\cdot}(\omega)-\int_{0}^{\cdot}G(s)\mathrm{d}\zeta_{s}(\omega)$.
By Lemma 3.5-(i), we have $y_{\cdot}$ has continuous trajectories under probability $R_{\ell}$, and hence $\tilde{R}_{\ell}$ is well-defined. On the other hand, for $R\in\mathcal{R}_{\rm opt}^{\epsilon}(\nu)$, Assumption 1-(A5) and (A6) yield that
$v^{*}+\epsilon$ $\geq\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}f(t,X_{t},R^{X}(t),u)q_{t}(\mathrm{d}u)\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}\right]$
$\geq\mathbb{E}^{R}\left[-\int_{0}^{T}M(1+|X_{t}|^{p}+M_{2}(\mu))\mathrm{d}t+M_{1}\int_{0}^{T}\int_{U}|u|^{p}q_{t}(\mathrm{d}u)\mathrm{d}t+c_{0}|\zeta_{T}|\right].$
Consequently, we have
$\sup_{R\in\mathcal{R}_{\rm opt}^{\epsilon}(\nu)}\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}|u|^{p}q_{t}(\mathrm{d}u)\mathrm{d}t+\left|\zeta_{T}\right|\right]<+\infty.$ (10)
Then, it follows from Corollary 2.9 in [20] that $(\tilde{R}_{\ell}|_{\mathcal{A}^{k}[0,T]})_{\ell\geq 1}$ is tight. Moreover, by Proposition B.3 in Lacker [24] and Kolmogorov’s tightness criterion, we can deduce that $(\tilde{R}_{\ell}|_{\mathcal{Q}\times C([0,T];\mathbb{R}^{n})})_{\ell\geq 1}$ is relatively compact in $\mathcal{P}_{2}(\mathcal{Q}\times C([0,T];\mathbb{R}^{n}))$.
We next prove that $(\tilde{R}_{\ell})_{\ell\geq 1}$ is tight. Since both $(\tilde{R}_{\ell}|_{\mathcal{Q}\times C([0,T];\mathbb{R}^{n})})_{\ell\geq 1}$ and $(\tilde{R}_{\ell}|_{\mathcal{A}^{k}[0,T]})_{\ell\geq 1}$ are tight, for any $\epsilon>0$, there exist some compact subsets $K\subset\mathcal{Q}$, $V\subset\mathcal{A}^{k}[0,T]$ and $C\subset C([0,T];\mathbb{R}^{n})$ such that $\sup_{\ell\geq 1}\tilde{R}_{\ell}(D^{n}[0,T]\times K\times V\times C)\geq 1-\epsilon$. Let $G(V):=\{\int_{0}^{\cdot}G(s)\mathrm{d}\zeta;~\zeta\in V\}$ and $D=C+G(V)=\left\{y+g;~y\in C,g\in G(V)\right\}$. Note that $C$ and $G(V)$ are compact in $C([0,T];\mathbb{R}^{n})$ and $\mathcal{A}^{n}[0,T]$, respectively. Then, they are both compact in $D^{n}[0,T]$ and so is $D$. Moreover, $(\tilde{R}_{\ell})_{\ell\geq 1}$ is supported on $\{x_{\cdot}=y_{\cdot}+\int_{0}^{\cdot}G(s)\mathrm{d}\zeta\}$, and thus $\sup_{\ell\geq 1}\tilde{R}_{\ell}(D\times K\times V\times C)=\sup_{\ell\geq 1}\tilde{R}_{\ell}(D^{n}[0,T]\times K\times V\times C)\geq 1-\epsilon$. Here, $D\times K\times V\times C$ is compact in $\Omega$, and hence $(R_{\ell})_{\ell\geq 1}$ is tight. Moreover, it follows from (10) that
$\sup_{R\in\mathcal{R}_{\rm opt}^{\epsilon}(\nu)}\mathbb{E}^{R}\left[\sup_{t\in[0,T]}\left|X_{t}\right|^{p}+\int_{0}^{T}\int_{U}|u|^{p}q_{t}(\mathrm{d}u)\mathrm{d}t+\sup_{t\in[0,T]}\left|Y_{t}\right|^{p}\right]<+\infty.$
[…]
$\tilde{R}\left(\left\{\tilde{\omega}\in\tilde{\Omega};~y_{\cdot}=x_{\cdot}+\int_{0}^{\cdot}G(s)\mathrm{d}\zeta_{s}\right\}\right)\geq\limsup_{\ell\to\infty}\tilde{R}_{\ell}\left(\left\{\tilde{\omega}\in\tilde{\Omega};~y_{\cdot}=x_{\cdot}+\int_{0}^{\cdot}G(s)\mathrm{d}\zeta_{s}\right\}\right)=1,$
which proves claim (i) of Lemma 3.5. For the martingale argument, because $(b,\sigma)$ are uniformly continuous in $\mu\in\mathcal{P}_{p}(\mathbb{R}^{n})$ (actually there exists a uniform Lipschitz constant) w.r.t. $(t,x,u)\in[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}$, for any $g\in C_{b}^{2}(\mathbb{R}^{n})$, it holds that
$\tilde{M}_{t}^{R_{\ell}}g(\omega)\to\tilde{M}_{t}^{R}g(\omega),\quad\ell\to\infty$ (11)
uniformly w.r.t. $(t,\omega)\in[0,T]\times\Omega$. Moreover, for any bounded $\mathcal{F}_{s}$-measurable random variable $\eta$ with $0\leq s<t\leq T$, we have
$\left|\mathbb{E}^{R}\left[(\tilde{M}_{t}^{R}g(\omega)-\tilde{M}_{s}^{R}g(\omega))\eta(\omega)\right]\right|\leq\mathbb{E}^{R_{\ell}}\left[\left|\left(\tilde{M}_{t}^{R}g(\omega)-\tilde{M}_{t}^{R_{n}}g(\omega)\right)\eta(\omega)\right|\right]$
$\quad+\mathbb{E}^{R_{\ell}}\left[\left|\left(\tilde{M}_{s}^{R}g(\omega)-\tilde{M}_{s}^{R_{\ell}}g(\omega)\right)\eta(\omega)\right|\right]+\left|\mathbb{E}^{R}\left[\tilde{M}_{t}^{R}g(\omega)\eta(\omega)\right]-\mathbb{E}^{R_{\ell}}\left[\tilde{M}_{t}^{R}g(\omega)\eta(\omega)\right]\right|$
$\quad+\left|\mathbb{E}^{R}\left[\tilde{M}_{s}^{R}g(\omega)\eta(\omega)\right]-\mathbb{E}^{R_{\ell}}\left[\tilde{M}_{s}^{R}g(\omega)\eta(\omega)\right]\right|+\left|\mathbb{E}^{R_{\ell}}\left[\left(\tilde{M}_{t}^{R_{\ell}}g(\omega)-\tilde{M}_{s}^{R_{\ell}}g(\omega)\right)\eta(\omega)\right]\right|$
$\quad=:I_{1}+I_{2}+I_{3}+I_{4}+I_{5},$
[…]
We are ready to present the main result of this section:
Theorem 3.9.
[…]
Proof.
Note that $v^{*}=\inf_{R\in\mathcal{R}_{\rm ad}(\nu)}\Gamma(R)=\inf_{R\in\mathcal{R}_{\rm opt}^{\epsilon}(\nu)}\Gamma(R)$. Then, it follows from Lemma 3.6 and Lemma 3.8 that $\mathcal{R}_{\rm opt}$ is nonempty. It also can be easily checked to be convex and closed. As a closed subset of the compact set $\mathcal{R}_{\rm opt}^{\epsilon}(\nu)$, it is hence compact. ∎
We also have the following corollary as a direct consequence of Lemma 3.4 and Theorem 3.9:
Corollary 3.10.
Under Assumption 1 and Assumption 4, there exists an optimal strict control $(\alpha,\zeta)$ to the constrained MFC problem (3).
4 SMP by the Lagrange Multipliers Method
The aim of this section is to characterize the optimal control by establishing the stochastic maximum principle (SMP) of the constrained MFC problem (1)-(3) with singular control.
Let us first introduce some basic spaces of functions and processes. Let $\mathcal{X}$ be the space $L^{2}([0,T];\mathbb{R}^{l})/\mathord{\sim}$ equipped with the inner product $\langle\gamma^{1},\gamma^{2}\rangle_{\mathcal{X}}=\int_{0}^{T}\langle\gamma_{t}^{1},\gamma_{t}^{2}\rangle\mathrm{d}t$. Then, $(\mathcal{X},\langle\cdot,\cdot\rangle_{\mathcal{X}})$ is a Hilbert space. Let $\mathbb{L}_{\mathbb{F}}^{2}$ be the set of $\mathbb{F}$-adapted processes $\beta=(\beta_{t})_{t\in[0,T]}$ taking values in $\mathbb{R}^{l}$ such that $\mathbb{E}[\int_{0}^{T}|\beta_{t}|^{2}\mathrm{d}t]<\infty$. As a result, $\mathbb{L}_{\mathbb{F}}^{2}$ is a Hilbert space equipped with the inner product $\langle\beta^{1},\beta^{2}\rangle_{\mathbb{L}_{\mathbb{F}}^{2}}=\mathbb{E}[\int_{0}^{T}\langle\beta^{1}_{t},\beta^{2}_{t}\rangle\mathrm{d}t]$ for $\beta^{1},\beta^{2}\in\mathbb{L}_{\mathbb{F}}^{2}$. It can be easily verified that $\mathscr{U}[0,T]$ is closed and convex in $(\mathbb{L}_{\mathbb{F}}^{2},\langle\cdot,\cdot\rangle_{\mathbb{L}_{\mathbb{F}}^{2}})$. Define $\mathbb{M}^{2}$ ($\subset\mathbb{L}_{\mathbb{F}}^{2}$) as the set of $\mathbb{F}$-adapted (continuous) $\mathbb{R}^{n}$-valued process $\eta=(\eta_{t})_{t\in[0,T]}$ satisfying the Itô’s representation that
$\eta_{t}=\eta_{0}+\int_{0}^{t}\eta^{1}_{s}\mathrm{d}s+\int_{0}^{t}\eta_{s}^{2}\mathrm{d}W_{s},$ (12)
where $\eta_{0}\in L^{2}(\Omega;\mathbb{R}^{n})$ is $\mathcal{F}_{0}$-measurable, $\eta^{1}=(\eta_{t}^{1})_{t\in[0,T]}$ and $\eta^{2}=(\eta_{t}^{2})_{t\in[0,T]}$ take values respectively in $\mathbb{R}^{n}$ and $\mathbb{R}^{n\times r}$, which are $\mathbb{F}$-adapted processes such that $\mathbb{E}[\int_{0}^{T}(|\eta_{t}^{1}|^{2}+|\eta_{t}^{2}|^{2})\mathrm{d}t]<\infty$. Thus, $\mathbb{M}^{2}$ is a Hilbert space under the inner product $\langle\cdot,\cdot\rangle_{\mathbb{M}^{2}}$ defined by
$\langle\eta,\delta\rangle_{\mathbb{M}^{2}}:=\mathbb{E}\left[\langle\eta_{0},\delta_{0}\rangle+\int_{0}^{T}\left(\langle\eta_{t}^{1},\delta_{t}^{1}\rangle+\mathrm{tr}\left(\eta_{t}^{2}(\delta_{t}^{2})^{\top}\right)\right)\mathrm{d}t\right],\quad\forall\eta,\delta\in\mathbb{M}^{2}.$ (13)
We use $\mathbb{L}_{\mathbb{F},\mathbb{R}}^{2}$ to represent the set of $\mathbb{F}$-adapted processes in $L^{2}([0,T]\times\Omega;\mathbb{R})/\mathord{\sim}$. Then, $\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}$ is a Hilbert space with the inner product $\langle\beta^{1},\beta^{2}\rangle_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}:=\mathbb{E}[\int_{0}^{T}\beta_{t}^{1}\beta_{t}^{2}\mathrm{d}t]$ for $\beta^{1},\beta^{2}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}$.
To adopt the methodology introduced in [6], we propose to embed the state process into a suitable Banach space. Without loss of generality, we can assume that for any singular control $\zeta\in\mathcal{A}^{k}[0,T]$, the associated radon measures $\mu_{i}$ for $i\in I$ of the trajectory $t\mapsto\zeta_{t}(\omega)$ is singular w.r.t. Lebesgue measure. Otherwise, we can decompose $\mu_{i}$ using the Lebesgue decomposition w.r.t. the Lebesgue measure and absorb the absolutely continuous (regular) term to the drift $b$. Consider the following space:
$\mathscr{X}:=\Big\{X=(X_{t})_{t\in[0,T]};~X\text{ is }\mathbb{F}\text{-adapted process with $X_{0}\in L^{2}(\Omega;\mathbb{R}^{n})$ and }$ (14)
$~X_{\cdot}=X_{0}+\int_{0}^{\cdot}X_{s}^{1}\mathrm{d}s+\int_{0}^{\cdot}X_{s}^{2}\mathrm{d}W_{s}+\int_{0}^{\cdot}G(s)\mathrm{d}X_{s}^{3}~\text{with}~(X^{1},X^{2},X^{3})\in\mathbb{L}_{\mathbb{F}}^{2}\times\mathbb{L}_{\mathbb{F}}^{2}\times\mathcal{A}^{k}[0,T]\Big\}.$
Consequently, $\mathscr{X}$ is a Banach space under the norm $\|\cdot\|_{\mathscr{X}}$ given by
$\|X\|_{\mathscr{X}}:=\|X^{1}\|_{\mathbb{L}_{\mathbb{F}}^{2}}+\|X^{2}\|_{\mathbb{L}_{\mathbb{F}}^{2}}+\sqrt{\mathbb{E}\left[\left|\int_{0}^{T}G(s)\mathrm{d}X_{s}^{3}\right|^{2}\right]},~~\forall X\in\mathscr{X}.$
By the assumption that, for any $X\in\mathscr{X}$, its 3rd component $X^{3}$ of the decomposition in $\mathscr{X}$ is singular w.r.t. to Lebesgue measure, one can conclude that $\mathscr{X}$ is isometric to the product space $\mathbb{M}^{2}\times\mathbb{G}^{2}$ under the product norm, where $(\mathbb{G}^{2},\|\cdot\|_{\mathbb{G}^{2}})$ is a Banach space defined by
$\mathbb{G}^{2}:=\left\{S=(S_{t})_{t\in[0,T]};~S_{\cdot}=\int_{0}^{\cdot}G(s)\mathrm{d}\eta_{s}~\text{for}~\eta=(\eta_{t})_{t\in[0,T]}\in\mathcal{A}^{k}[0,T]\right\},$
$\|S\|_{\mathbb{G}^{2}}:=\sqrt{\mathbb{E}\left[\left|\int_{0}^{T}G(s)\mathrm{d}\eta_{s}\right|^{2}\right]}.$
Since $\int_{0}^{\cdot}G(s)\mathrm{d}\eta_{s}\in(\mathcal{M}[0,T])^{n}$ for $\eta\in\mathcal{A}^{k}[0,T]$, we can identify $\mathbb{G}^{2}$ as a subset of $L^{2}(\Omega;(\mathcal{M}[0,T])^{n})$ with $(\mathcal{M}[0,T])^{n}$ equipped with the norm $\|\mu\|_{(\mathcal{M}[0,T])^{n}}=\sqrt{\sum_{i=1}^{n}\|\mu_{i}\|^{2}_{\mathrm{TV}}}$. Thus, $\mathscr{X}$ can be viewed as a subset of $\mathbb{T}^{2}:=\mathbb{M}^{2}\times L^{2}(\Omega;(\mathbb{M}[0,T])^{n})$ under the product norm. For any $\lambda=(\lambda^{1},\lambda^{2})\in\mathbb{T}^{2,*}\simeq\mathbb{M}^{2}\times L^{2}(\Omega;(C([0,T];\mathbb{R})^{**})^{n})$, the dual pair of $X=L+S\in\mathscr{X}$ for $(L,S)\in\mathbb{M}^{2}\times\mathbb{G}^{2}$ is defined by
$\langle X,\lambda\rangle_{\mathbb{T}^{2},\mathbb{T}^{2,*}}=\langle L,\lambda^{1}\rangle_{\mathbb{M}^{2}}+\langle S,\lambda^{2}\rangle_{L^{2}(\Omega;(\mathbb{M}[0,T])^{k}),L^{2}(\Omega;(C([0,T];\mathbb{R})^{**})^{k})}.$
Following Lemma 3.6 in [6], let us introduce
$\mathtt{X}:=\mathbb{T}^{2}\times\mathbb{L}_{\mathbb{F}}^{2}\times L^{2}(\Omega;(\mathbb{M}[0,T])^{n}),\quad\mathtt{Y}_{i}:=\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}},~\forall i\in I,\quad\mathtt{H}:=\mathbb{T}^{2}.$ (15)
The norm for $\mathtt{X}$ is defined to be the (default) product norm and we take $\|\cdot\|_{\mathtt{Y}_{i}}$ $(i\in I)$ as the norm $\|\cdot\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}$. Moreover, define $J:\mathtt{X}\to\mathbb{R}$, $\Phi_{i}:\mathtt{X}\to\mathtt{Y}_{i}$ $(i\in I)$ and $F:\mathtt{X}\to\mathtt{H}$ by, for $(X,\alpha,\zeta)\in\mathtt{X}$,
$J(X,\alpha,\zeta)$ $:=\mathbb{E}\left[\int_{0}^{T}f(t,X_{t},\mathcal{L}(X_{t}),\alpha_{t})\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}\right],$
$\Phi_{i}(X,\alpha,\zeta)(t)$ $:=-\phi^{i}(t,X_{t},\mathcal{L}(X_{t}),\alpha_{t}),~~~\forall t\in[0,T],~i\in I,$ (16)
$F(X,\alpha,\zeta)(\cdot)$ $:=X_{\cdot}-\xi-\int_{0}^{\cdot}b(s,X_{s},\mathcal{L}(X_{s}),\alpha_{s})\mathrm{d}s-\int_{0}^{\cdot}\sigma(s,X_{s},\mathcal{L}(X_{s}),\alpha_{s})\mathrm{d}W_{s}-\int_{0}^{\cdot}G(s)\mathrm{d}\zeta_{s}.$
The well-posedness of $\Phi_{i}$ is guaranteed by Assumption 1-(A2). Thus, we can formulate the following optimization problem (equivalent to the constrained MFC problem (3)) with constraints that
$\begin{cases}\displaystyle\inf~J(X,\alpha,\zeta)\\[3.99994pt] \displaystyle\mathrm{s.t.}~~(X,\alpha,\zeta)\in C:=\mathscr{X}\times\mathscr{U}[0,T]\times\mathcal{A}^{k}[0,T],\\[3.99994pt] \displaystyle~~~~~~~\Phi_{i}(X,\alpha,\zeta)\leq_{K}0,~~\forall i\in I,\\[3.99994pt] \displaystyle\displaystyle~~~~~~~F(X,\alpha,\zeta)=0,\end{cases}$ (17)
where $C$ in (17) is closed and convex in $\mathtt{X}$, and $K$ is a closed and convex cone given by
$\displaystyle K:=\{f\in\mathbb{L}_{\mathbb{F},\mathbb{R}}^{2};~f(t,\omega)\geq 0,~\forall(t,\omega)\in[0,T]\times\Omega\}=:\mathbb{L}_{\mathbb{F},\mathbb{R}+}^{2}.$ (18)
By using Lemma 3.11 in [6], the FJ optimality condition holds for the constrained optimization problem (17) in the following sense.
Lemma 4.1 (FJ optimality condition for problem (17)).
Let $(\widehat{X},\widehat{\alpha},\widehat{\zeta})\in C$ be an optimal pair to problem (17). Then, there exist multipliers $r_{0}\geq 0$, $\eta^{i}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $i\in I$ and $\lambda=(\lambda_{1},\lambda_{2})\in\mathbb{M}^{2}\times L^{2}(\Omega;(C([0,T],\mathbb{R})^{**})^{n})$, which are not all zeros, such that
$\mathbb{E}\left[\int_{0}^{T}\Phi_{i}(\widehat{X},\widehat{\alpha},\widehat{\zeta})(t)\eta^{i}_{t}\mathrm{d}t\right]=0,~~\forall i\in I,$
$r_{0}+\sum_{i\in I}\|\eta^{i}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}+\|\lambda\|_{\mathbb{M}^{2}\times L^{2}(\Omega;(C([0,T],\mathbb{R})^{**})^{n})}=1,$
$-\xi_{1}=r_{0}D_{X}J(\widehat{X},\widehat{\alpha},\widehat{\zeta})(\cdot)+\sum_{i\in I}\mathbb{E}\left[\int_{0}^{T}D_{X}\Phi_{i}(\widehat{X},\widehat{\alpha},\widehat{\zeta})(\cdot)(t)\eta^{i}_{t}\mathrm{d}t\right]$
$\qquad\quad-\left\langle\lambda,D_{X}F(\widehat{X},\widehat{\alpha},\widehat{\zeta})(\cdot)\right\rangle_{\mathtt{H}^{*},\mathtt{H}}~~\text{\rm in}~\mathbb{T}^{2},$
$-\xi_{2}=r_{0}D_{\alpha}J(\widehat{X},\widehat{\alpha},\widehat{\zeta})(\cdot)+\sum_{i\in I}\mathbb{E}\left[\int_{0}^{T}D_{\alpha}\Phi_{i}(\widehat{X},\widehat{\alpha},\widehat{\zeta})(\cdot)(t)\eta^{i}_{t}\mathrm{d}t\right]$
$\qquad\quad-\left\langle\lambda,D_{\alpha}F(\widehat{X},\widehat{\alpha},\widehat{\zeta})(\cdot)\right\rangle_{\mathtt{H}^{*},\mathtt{H}}~~\text{\rm in}~\mathbb{L}_{\mathbb{F}}^{2},$
$-\xi_{3}=r_{0}D_{\zeta}J(\widehat{X},\widehat{\alpha},\widehat{\zeta})(\cdot)+\sum_{i\in I}\mathbb{E}\left[\int_{0}^{T}D_{\zeta}\Phi_{i}(\widehat{X},\widehat{\alpha},\widehat{\zeta})(\cdot)(t)\eta^{i}_{t}\mathrm{d}t\right]$
$\quad\qquad-\left\langle\lambda,D_{\zeta}F(\widehat{X},\widehat{\alpha},\widehat{\zeta})(\cdot)\right\rangle_{\mathtt{H}^{*},\mathtt{H}}~~\text{\rm in}~L^{2}(\Omega;(C([0,T];\mathbb{R})^{**})^{n}),$
where $\xi=(\xi_{1},\xi_{2},\xi_{3})\in N_{C}(\widehat{X},\widehat{\alpha},\widehat{\zeta})\subset\mathtt{X}^{*}$ and $\|\cdot\|_{\rm TV}$ is the total variation norm on the space of (signed) Radon measures.
Then, we have the following main result in this section.
Theorem 4.2 (SMP for constrained MFC problem (3) with singular control ).
Let $(\widehat{X},\widehat{\alpha},\widehat{\zeta})\in C$ be an optimal solution to problem (3). There exist some $r_{0}\geq 0$ and $\eta^{i}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $i\in I$ such that
  • (i)
    for $i\in I$, $\eta_{t}^{i}\geq 0$, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-$\mathrm{a.s.}$ and $\eta^{i}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$, let $A^{j}=(A_{t}^{i})_{t\in[0,T]}$ be the increasing process that $A_{t}^{i}=\int_{0}^{t}\eta_{s}^{i}\mathrm{d}s$. Then, for $\mathbb{P}$-$\mathrm{a.s.}$ $\omega$, the process $A^{i}(\omega)$ only increases on the set $\{t\in[0,T];~\phi^{j}(t,\widehat{X}_{t}(\omega),\widehat{\mu}_{t},\widehat{\alpha}_{t}(\omega))=0\}$, $\mathrm{d}t$-$\mathrm{a.s.}$, with $\widehat{\mu}_{t}=\mathcal{L}(\widehat{X}_{t})$;
  • (ii)
    it holds that $r_{0}+\sum_{i\in I}\|\eta^{i}\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}=1$;
  • (iii)
    $\mathbb{P}\big(r_{0}c_{i}(t)+G_{i}(t)^{\top}Y_{t}\geq 0,~\forall(t,i)\in[0,T]\times\{1,\ldots,k\}\big)=1$ with $c_{i}(t)$ and $G_{i}(t)$ being the $i$-th row of $c(t)$ and $G(t)$, respectively. It also holds that
    $\mathbb{P}\left(\sum_{i=1}^{k}\mathbf{1}_{\left\{r_{0}c_{i}(t)+G_{i}(t)^{\top}Y_{t}>0\right\}}\mathrm{d}\widehat{\zeta}_{t}^{i}=0\right)=1.$ (19)
Furthermore, the stochastic (first-order) minimum condition holds that, $\mathrm{d}t\otimes\mathrm{d}\mathbb{P}$-a.s.
$\left\langle\nabla_{u}H^{r_{0}}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t},Y_{t},Z_{t})-\sum_{i\in I}\nabla_{u}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})\eta_{t}^{i},u-\widehat{\alpha}_{t}\right\rangle\geq 0,\quad\forall u\in U.$ (20)
The Hamiltonian $H^{r_{0}}:[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n\times r}\mapsto\mathbb{R}$ is defined by
$H^{r_{0}}(t,x,u,\mu,y,z):=\langle b(t,x,\mu,u),y\rangle+\mathrm{tr}\left(\sigma(t,x,\mu,u)z^{\top}\right)+r_{0}f(t,x,\mu,u),$ (21)
and the process pair $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ taking values in $\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ is the unique solution to the constrained BSDE:
$Y_{t}$ $=\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime},Y_{s}^{\prime},Z_{s}^{\prime})(\widehat{X}_{s})\right]\mathrm{d}s$
$\quad-\sum_{i\in I}\int_{t}^{T}\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})\eta^{i}_{s}\mathrm{d}s-\sum_{i\in I}\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s}){\eta^{i}_{s}}^{\prime}\right]\mathrm{d}s$
$\quad-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}.$
Here, $(\widehat{X}^{\prime},\widehat{\alpha}^{\prime},Y^{\prime},Z^{\prime},\eta^{i^{\prime}})$ is an independent copy of $(\widehat{X},\widehat{\alpha},Y,Z,\eta^{i})$ defined on the probability space $(\Omega^{\prime},\mathcal{F}^{\prime},P^{\prime})$.
Proof.
[…]
$\langle\xi_{1},\widehat{X}-X\rangle_{\mathbb{T}^{2}}+\langle\xi_{2},\widehat{\alpha}-\alpha\rangle_{\mathbb{L}_{\mathbb{F}}^{2}}+\langle\xi_{3},\zeta-\widehat{\zeta}\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathbb{M}[0,T])^{n}}\geq 0.$
Taking $\alpha=\widehat{\alpha}$ and $\zeta=\widehat{\zeta}$, we have $\xi_{1}=0$, $\langle\xi_{2},\widehat{\alpha}-\alpha\rangle_{\mathbb{L}_{\mathbb{F}}^{2}}\geq 0$ and $\langle\xi_{3},\zeta-\widehat{\zeta}\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathbb{M}[0,T])^{n}}$ $\geq 0$ similar as [6] (there is a slight change in the proof of Lemma 3.10 to fit into our setting here). Then, we can conclude from the structure of Lions derivative (c.f. Section 5.2.1 in [10]) and Lemma 4.1 that
$0=r_{0}\mathbb{E}\left[\int_{0}^{T}\langle\nabla_{x}f(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t}),L_{t}\rangle\mathrm{d}t+\mathbb{E}^{\prime}\left[\int_{0}^{T}\langle\partial_{\mu}f(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime}),L_{t}^{\prime}\rangle\mathrm{d}t\right]\right]$
$-\sum_{i\in I}\mathbb{E}\left[\int_{0}^{T}\left\langle\nabla_{x}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})+\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime})\right],L_{t}^{\prime}\right\rangle\eta^{i}_{t}\mathrm{d}t\right]$
$-\mathbb{E}\left[\langle\lambda_{1}(0),L_{0}\rangle+\int_{0}^{T}\langle\lambda_{1}^{1}(t),L_{t}^{1}-\nabla_{x}b(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})L_{t}\rangle\mathrm{d}t+\int_{0}^{T}\mathrm{tr}\left(\lambda_{1}^{2}(t)(L_{t}^{2}-\nabla_{x}\sigma(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})L_{t})^{\top}\right)\mathrm{d}t\right$.
$+\left.\int_{0}^{T}\left\langle\lambda_{1}^{1}(t),\mathbb{E}^{\prime}\left[\partial_{\mu}b(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime})L_{t}^{\prime}\right]\right\rangle\mathrm{d}t+\int_{0}^{T}\mathrm{tr}\left(\lambda_{1}^{2}(t)\left(\mathbb{E}^{\prime}\left[\partial_{\mu}\sigma(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime})L_{t}^{\prime}\right]\right)^{\top}\right)\mathrm{d}t\right]$
$\qquad+\left\langle\lambda_{2},\nu\right\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathcal{M}[0,T])^{n}},$ (22)
where $\lambda_{1}\in\mathbb{M}^{2}$ has the decomposition $\lambda_{1}(\cdot)=\lambda_{1}(0)+\int_{0}^{\cdot}\lambda_{1}^{1}(s)\mathrm{d}s+\int_{0}^{\cdot}\lambda_{1}^{2}(s)\mathrm{d}W_{s}$, $L\in\mathbb{T}^{2}$ has the representation $L_{\cdot}=L_{0}+\int_{0}^{\cdot}L_{s}^{1}\mathrm{d}s+\int_{0}^{\cdot}L_{s}^{2}\mathrm{d}W_{s}+\int_{0}^{\cdot}G(s)\mathrm{d}L_{s}^{3}$, and $\nu$ is the associated radon measure of the mapping $\int_{0}^{\cdot}G(s)\mathrm{d}L_{s}^{3}\in\mathrm{BV}^{n}[0,T]$. Define an element $p=(p_{t})_{t\in[0,T]}$ in $\mathbb{M}^{2}$ that, for $t\in[0,T]$,
$p_{t}$ $=-\sum_{i\in I}\int_{t}^{T}\left(\nabla_{x}\phi^{i}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})+\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s}){\eta^{i}_{s}}^{\prime}\right]\right)\mathrm{d}s$
$\quad+\int_{t}^{T}\left[r_{0}\nabla_{x}f(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})+\nabla_{x}b(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})\lambda_{1}^{1}(s)+\nabla_{x}\sigma(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})\lambda_{1}^{2}(s)\right]\mathrm{d}s$
$\quad+\int_{t}^{T}\mathbb{E}^{\prime}\left[r_{0}\partial_{\mu}f(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s})+\partial_{\mu}b(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s}){\lambda_{1}^{1}(s)}^{\prime}\right$.
$\qquad\qquad+\left.\partial_{\mu}\sigma(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s}){\lambda_{1}^{2}(s)}^{\prime}\right]\mathrm{d}s-\int_{t}^{T}\lambda_{1}^{2}(s)\mathrm{d}W_{s},$ (23)
where $MN:=\sum_{i=1}^{r}M_{i}L_{i}\in\mathbb{R}^{n}$ for any $M=(M_{1},\ldots,M_{r})\in\mathbb{R}^{n\times n\times r}$ and $L=(L_{1},\ldots,L_{r})\in\mathbb{R}^{n\times r}$. By construction of the above $L\in\mathbb{T}^{2}$, we have
$\langle\lambda_{1}^{1}(t),L_{t}^{1}\rangle\mathrm{d}t=\langle\lambda_{1}^{1}(t)-p(t),\mathrm{d}L_{t}\rangle+\langle p(t),\mathrm{d}L_{t}\rangle-\langle\lambda_{1}^{1},\nu\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathcal{M}[0,T])^{n}}.$ (24)
By using (4) together with Itô’s rule, we can conclude by Fubini’s theorem that
$\mathbb{E}\left[\int_{0}^{T}\langle p_{t},\mathrm{d}L_{t}\rangle\right]=-\mathbb{E}\left[\int_{0}^{T}\langle L_{t},\mathrm{d}p_{t}\rangle\mathrm{d}t\right]-\mathbb{E}\left[\langle p,L\rangle_{T}\right]$
$=\mathbb{E}\left[\int_{0}^{T}\left\langle r_{0}\nabla_{x}f(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t}),L_{t}\right\rangle\mathrm{d}t+\int_{0}^{T}\left\langle r_{0}\mathbb{E}^{\prime}\left[\partial_{\mu}f(t,\widehat{X}_{t}^{\prime},\widehat{\mu}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t})\right],L_{t}\right\rangle\mathrm{d}t\right]$
$+\mathbb{E}\left[\int_{0}^{T}\left(\langle\lambda_{1}^{1}(t),\nabla_{x}b(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})L_{t}\rangle+\mathrm{tr}\left(\lambda_{1}^{2}(t)\left(\nabla_{x}\sigma(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})L_{t}\right)^{\top}\right)\right)\right]$
$+\mathbb{E}\left[\int_{0}^{T}\left\langle\lambda_{1}^{1}(t),\mathbb{E}^{\prime}\left[\partial_{\mu}b(t,\widehat{X}_{t}^{\prime},\widehat{\mu}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t})L_{t}\right]\right\rangle\mathrm{d}t-\langle p_{0},L_{0}\rangle\right]-\mathbb{E}\left[\mathrm{tr}\left(\lambda_{1}^{2}(t)L_{t}^{2}\right)\mathrm{d}t\right]$
$+\mathbb{E}\left[\int_{0}^{T}\mathrm{tr}\left(\lambda_{1}^{2}(t)\left(\mathbb{E}^{\prime}\left[\partial_{\mu}\sigma(t,\widehat{X}_{t}^{\prime},\widehat{\mu}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t})L_{t}\right]\right)^{\top}\right)\mathrm{d}t\right]$
$-\sum_{i\in I}\mathbb{E}\left[\int_{0}^{T}\left\langle\nabla_{x}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t}),L_{t}\right\rangle\mathrm{d}t+\int_{0}^{T}\left\langle\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(t,\widehat{X}_{t}^{\prime},\widehat{\mu}_{t},\widehat{\alpha}_{t}^{\prime})(\widehat{X}_{t})\right],L_{t}\right\rangle\mathrm{d}t\right].$
Inserting (24) together with the above equality into (4), we derive
$\mathbb{E}\left[\langle p_{0}-\lambda_{1}(0),L_{0}\rangle+\int_{0}^{T}\langle p_{t}-\lambda_{1}^{1}(t),\mathrm{d}L_{t}\rangle+\langle\lambda_{2}-\lambda_{1}^{1},\nu\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathcal{M}[0,T])^{n}}\right]=0.$
We thus deduce that $p_{0}=\lambda_{1}(0)$ and $p$ is indistinguishable from $\lambda_{1}^{1}$ by the arbitrariness of $L$. As a result, $\lambda_{1}^{1}$ solves (5.3). Moreover, we have $G^{\top}\lambda_{2}=G^{\top}\lambda_{1}^{1}$ in $(C([0,T];\mathbb{R})^{**})^{k}$, and hence $G^{\top}\lambda_{2}\in C([0,T];\mathbb{R}^{k})$. Here, $G$ can be viewed a mapping in $\mathcal{L}((C([0,T];\mathbb{R})^{**})^{n};$ $((C([0,T];\mathbb{R})^{**})^{k}))$. By the optimality condition for $\widehat{\alpha}\in\mathscr{U}[0,T]$, we have
$0$ $\leq\mathbb{E}\left[\int_{0}^{T}\left\langle r_{0}\nabla_{u}f(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})+\nabla_{u}b(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})^{\top}\lambda_{t}^{1}+\nabla_{u}\sigma(t,\widehat{X}_{t},\widehat{\rho}_{t},\widehat{\alpha}_{t})^{\top}\lambda_{t}^{2},\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t\right$.
$\quad-\left.\sum_{i\in I}\int_{0}^{T}\left\langle\nabla_{u}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})\eta_{t}^{i},\alpha_{t}-\widehat{\alpha}_{t}\right\rangle\mathrm{d}t\right].$ (25)
Exploiting the fact that $(\lambda^{1}_{1},\lambda_{1}^{2})$ solves (5.3), we then can derive that (4) indicates (20) by the arbitrariness of $\alpha\in\mathscr{U}[0,T]$. By the optimality condition for $\widehat{\zeta}\in\mathcal{A}^{k}[0,T]$, we arrive at
$0$ $\leq-\mathbb{E}\left[\langle\xi_{3},\eta-\widehat{\zeta}\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathbb{M}[0,T])^{n}}\right]$
$=\mathbb{E}\left[r_{0}\int_{0}^{T}c(t)\mathrm{d}(\eta_{t}-\hat{\zeta}_{t})+\left\langle\lambda_{2},\int_{0}^{T}G(t)\mathrm{d}(\eta_{t}-\hat{\zeta}_{t})\right\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathbb{M}[0,T])^{n}}\right]$
$=\mathbb{E}\left[r_{0}\int_{0}^{T}c(t)\mathrm{d}(\eta_{t}-\hat{\zeta}_{t})+\int_{0}^{T}G(t)^{\top}\lambda^{1}_{1}(t)\mathrm{d}(\eta_{t}-\hat{\zeta}_{t})\right].$
For the last equality, we have used the fact that $G^{\top}\lambda_{2}=G^{\top}\lambda_{1}^{1}$. Taking $\eta_{t}=\hat{\zeta}_{t}+\delta_{t}\in\mathcal{A}^{k}[0,T]$ for any $\delta=(\delta_{t})_{t\in[0,T]}\in\mathcal{A}^{k}[0,T]$, we derive that
$\mathbb{P}\left(r_{0}c_{i}(t)+G_{i}(t)^{\top}Y_{t}\geq 0,~\forall(t,i)\in[0,T]\times\{1,\ldots,k\}\right)=1.$
Moreover, we take $\eta_{t}\equiv 0$, this results in (19). Thus, the proof of the theorem is complete by just noting that (i) follows from Lemma 4.1. ∎
Remark 4.3.
When there are no constraints (i.e. $m=0$) and distribution-dependent terms, Theorem 4.2 together with the constraints qualification (CQ) condition derived in [6] will reduce to the SMP as in [3].
5 Uniqueness and Stability of Solution to Constrained FBSDE
The aim of this section is to study the uniqueness and stability of solutions to the following generalized type of constrained FBSDE derived in Theorem 4.2. Without loss of generality, we assume that $d=1$:
$\widehat{X}_{t}$ $=\xi+\int_{0}^{t}b(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})\mathrm{d}W_{s}+\int_{0}^{t}G(s)\mathrm{d}\widehat{\zeta}_{s},$
$Y_{t}$ $=\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\mu},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H^{r_{0}}(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime},Y_{s}^{\prime},Z_{s}^{\prime})(\widehat{X}_{s})\right]\mathrm{d}s$
$\quad-\int_{t}^{T}\nabla_{x}\phi(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})\eta_{s}\mathrm{d}s-\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}\phi(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime})(\widehat{X}_{s}){\eta_{s}}^{\prime}\right]\mathrm{d}s$
$\quad-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}$ (26)
subject to $|r_{0}|+\|\eta\|_{\mathbb{L}_{\mathbb{F},\mathbb{R}}}=1$, $\widehat{\mu}_{t}=\mathcal{L}(\widehat{X}_{t})$, $\mathbb{P}\left(G(t)Y_{t}+c(t)\geq 0,~\forall t\in[0,T]\right)=1$, and $\mathrm{d}t\otimes\mathrm{d}\mathbb{P}$-a.s.
$\begin{cases}\displaystyle\left\langle\nabla_{u}H^{r_{0}}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t},Y_{t},Z_{t})-\sum_{i\in I}\nabla_{u}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})\eta_{t}^{i},u-\widehat{\alpha}_{t}\right\rangle\geq 0,~\forall u\in U,\\[10.00002pt] \displaystyle\int_{0}^{T}\phi(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})\eta_{t}\mathrm{d}t=0,\quad\phi\big(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t}\big)\geq 0,\quad\eta_{t}\geq 0,\\[10.00002pt] \displaystyle~\mathbb{P}\left(\sum_{i=1}^{k}\mathbf{1}_{\{r_{0}c_{i}(t)+G_{i}(t)^{\top}Y_{t}>0\}}\mathrm{d}\widehat{\zeta}_{t}^{i}=0\right)=1,\end{cases}$ (27)
which characterizes the necessary optimality condition of the constrained MFC problem (3) with singular control. To prove the uniqueness of solution $(\widehat{X},\widehat{\alpha},\widehat{\zeta},Y,Z,\eta)=(\widehat{X}_{t},\widehat{\alpha}_{t},\widehat{\zeta}_{t},Y_{t},Z_{t},\eta_{t})_{t\in[0,T]}$ satisfying (5)-(27), we first give the definition of uniqueness of solutions to Eq. (5)-(27):
Definition 5.1 (Uniqueness of solutions to FBSDE (5)-(27)).
Let $(\widehat{X}^{i},\widehat{\alpha}^{i},\widehat{\zeta}^{i},Y^{i},Z^{i},\eta^{i})$ $(\widehat{X}_{t}^{i},\widehat{\alpha}_{t}^{i},\widehat{\zeta}_{t}^{i},Y_{t}^{i},Z_{t}^{i},\eta_{t}^{i})_{t\in[0,T]}$ for $i=1,2$ be $\mathbb{F}$-adapted processes valued in $\mathbb{R}^{n}\times U\times\mathbb{R}_{+}^{k}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\times\mathbb{R}_{+}$ satisfying the coupling constrained FBSDE (5)-(27). The solution of FBSDE (5)-(27) is said to be unique if $\widehat{X}^{1}=\widehat{X}^{2}$, $\widehat{\alpha}^{1}=\widehat{\alpha}^{2}$, $Y^{1}=Y^{2}$, $Z^{1}=Z^{2}$, $\eta^{1}=\eta^{2}$ and $\int_{0}^{\cdot}G(s)\mathrm{d}\widehat{\zeta}_{s}^{1}=\int_{0}^{\cdot}G(s)\mathrm{d}\widehat{\zeta}_{s}^{2}$.
Remark 5.2.
In Definition 5.1, we only require $\int_{0}^{\cdot}G(s)\mathrm{d}\widehat{\zeta}_{s}^{1}=\int_{0}^{\cdot}G(s)\mathrm{d}\widehat{\zeta}_{s}^{2}$ rather than $\widehat{\zeta}^{1}=\widehat{\zeta}^{2}$ . Once the existence of optimal triple $(\widehat{X},\widehat{\alpha},\widehat{\zeta})=(\widehat{X}_{t},\widehat{\alpha}_{t},\widehat{\zeta}_{t})_{t\in[0,T]}$ is guaranteed, the FBSDE (5)-(27) has a solution with the backward component $(Y,Z,\eta)=(Y_{t},Z_{t},\eta_{t})_{t\in[0,T]}$ being the multipliers given in the last section. Then, by using Corollary 3.10, we can conclude that, under Assumption 1 and Assumption 4, there exists a solution to the generalized type constrained FBSDE (5)-(27).
We now verify the CQ condition as in (1.4) of [33] to establish the KKT optimality condition (i.e., we verify $r_{0}=1$ in Lemma 4.1).
Lemma 5.3.
Under Assumption 3-(C1), the KKT optimality condition holds. More precisely, if $(\widehat{X},\widehat{\alpha},\widehat{\zeta})\in C:=\mathbb{M}^{2}\times\mathscr{U}[0,T]\times\mathcal{A}^{k}[0,T]$ is an optimal solution to the constrained MFC problem (3), there exists some $\eta\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ such that
  • (i)
    $\mathbb{E}\left[\int_{0}^{T}\phi(t,\widehat{X}_{t},\mathcal{L}(\widehat{X}_{t}))\eta_{t}\mathrm{d}t\right]=0$;
  • (ii)
    $\mathbb{P}(c_{i}(t)+G_{i}(t)^{\top}Y_{t}\geq 0,~\forall(i,t)\in\{1,\ldots,k\}\times[0,T])=1$, where $c_{i}(t)$ and $G_{i}(t)$ are the $i$-th row of $c(t)$ and $G(t)$, respectively. It also holds that
    $\mathbb{P}\left(\sum_{i=1}^{k}\mathbf{1}_{\{c_{i}(t)+G_{i}(t)^{\top}Y_{t}>0\}}\mathrm{d}\widehat{\zeta}_{t}^{i}=0\right)=1.$ (28)
Moreover, the stochastic (first-order) minimum condition holds that, $\mathrm{d}t\otimes\mathrm{d}\mathbb{P}$-a.s.
$\left\langle\nabla_{u}H(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t},Y_{t},Z_{t}),u-\widehat{\alpha}_{t}\right\rangle\geq 0,\quad\forall u\in U.$ (29)
Here, the Hamiltonian $H(t,x,\mu,u,y,z)=H^{1}(t,x,\mu,u,y,z)$ for $(t,x,\mu,u,y,z)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$, and the process pair $(Y,Z)=(Y_{t},Z_{t})_{t\in[0,T]}$ taking values in $\mathbb{R}^{n}\times\mathbb{R}^{n\times r}$ is the unique solution to the following general type of constrained BSDE:
$Y_{t}$ $=\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime},Y_{s}^{\prime},Z_{s}^{\prime})(\widehat{X}_{s})\right]\mathrm{d}s$
$\quad-\int_{t}^{T}\nabla_{x}\phi(s,\widehat{X}_{s},\widehat{\mu}_{s})\eta_{s}\mathrm{d}s-\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}\phi(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s})(\widehat{X}_{s})\eta_{s}^{\prime}\right]\mathrm{d}s-\int_{t}^{T}Z_{s}\mathrm{d}W_{s}.$ (30)
Proof.
This is a direct consequence of Lemma 3.9 in [6] and Theorem 3.1 of [33]. ∎
The strict convexity imposed on the Hamiltonian $H$ w.r.t. the control variable $u\in U$ given in (C3) of Assumption 3 yields the result that once $\langle\nabla_{u}H(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t},Y_{t},Z_{t}),u-\widehat{\alpha}_{t}\rangle\geq 0$ for all $u\in U$, we must have $\widehat{\alpha}_{t}=\hat{u}(t,\widehat{X}_{t},\widehat{\mu}_{t},Y_{t},Z_{t})$. Then, we have
Theorem 5.4 (Uniqueness of solutions to FBSDE).
Let Assumption 2 and Assumption 3 hold. The constrained FBSDE (5)-(27) has a unique solution in the sense of Definition 5.1.
Proof.
Let $(X^{i}_{t},\zeta^{i}_{t},Y^{i}_{t},Z^{i}_{t},\eta_{t}^{i})_{t\in[0,T]}$ ($i=1,2$) be two solutions to the FBSDE (5)-(27) satisfying $X^{1}_{0}=X_{0}^{2}$ and $Y_{T}^{1}=Y_{T}^{2}$. Set $\bar{X}_{t}:=X_{t}^{1}-X_{t}^{2}$, $\bar{\zeta}_{t}:=\zeta_{t}^{1}-\zeta_{t}^{2}$, $\bar{Y}_{t}:=Y_{t}^{1}-Y_{t}^{2}$, $\bar{Z}_{t}:=Z_{t}^{1}-Z_{t}^{2}$ and $\bar{\eta}_{t}:=\eta_{t}^{1}-\eta_{t}^{2}$ for $t\in[0,T]$. Applying Itô’s formula to $\bar{X}_{t}\bar{Y}_{t}$, we arrive at
$\bar{X}_{T}\bar{Y}_{T}$ $=\int_{0}^{T}\left\{-\Delta\bar{F}(t)\bar{X}_{t}+\bar{Y}\Delta\bar{b}(t)+\mathrm{tr}(\bar{Z}_{t}\Delta\bar{\sigma}(t))\right\}\mathrm{d}t+\int_{0}^{T}\left(\bar{X}_{t}\bar{Z}_{t}+\bar{Y}_{t}\Delta\sigma(t)\right)\mathrm{d}W_{t}$
$\quad+\int_{0}^{T}\left(A(t)\bar{\eta}_{t}+B(t)\mathbb{E}[\bar{\eta}_{t}]\right)\bar{X}_{t}\mathrm{d}t+\int_{0}^{T}G(t)\bar{Y}_{t}\mathrm{d}\bar{\zeta}_{t},$
where $\Delta\bar{F}(t):=\bar{F}(t,X_{t}^{1},\mathcal{L}(X_{t}^{1}),Y_{t}^{1},Z_{t}^{1})-\bar{F}(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2})$, $\Delta\bar{b}(t):=\bar{b}(t,X_{t}^{1},\mathcal{L}(X_{t}^{1}),Y_{t}^{1},Z_{t}^{1})-\bar{b}(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2})$ and $\Delta\bar{\sigma}(t):=\bar{\sigma}(t,X_{t}^{1},\mathcal{L}(X_{t}^{1}),Y_{t}^{1},Z_{t}^{1})-\bar{\sigma}(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2})$. Taking expectations on both sides and noticing $\bar{Y}_{T}=0$, we get
[…]
$\qquad~+\left.(\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{1}),Y_{t}^{2},Z_{t}^{2})-\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2}))(X_{t}^{1}-X_{t}^{2},Y_{t}^{1}-Y_{t}^{2},Z_{t}^{1}-Z_{t}^{2})\right\}\mathrm{d}t$
$\qquad~+\left.\int_{0}^{T}\bar{\eta}_{t}\left\{a(t)\bar{X}_{t}+b(t)\mathbb{E}[\bar{X}_{t}]\right\}\mathrm{d}t+\int_{0}^{T}G(t)\bar{Y}_{t}\mathrm{d}\bar{\zeta}_{t}\right].$
Here, we applied Fubini’s theorem in the last line. Note that, it holds that
$\mathbb{E}\left[\int_{0}^{T}\bar{\eta}_{t}\left\{A(t)\bar{X}_{t}+B(t)\mathbb{E}[\bar{X}_{t}]\right\}\mathrm{d}t\right]=\mathbb{E}\left[\int_{0}^{T}\bar{\eta}_{t}\left\{\phi(t,X_{t}^{1},\mathcal{L}(X_{t}^{1}))-\phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}))\right\}\mathrm{d}t\right]$
$\qquad\qquad=-\mathbb{E}\left[\int_{0}^{T}\left\{\phi(t,X_{t}^{1},\mathcal{L}(X_{t}^{1}))\eta_{t}^{2}+\phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}))\eta_{t}^{1}\right\}\mathrm{d}t\right]\leq 0,$
where we used the fact that $\mathbb{E}[\int_{0}^{T}\phi(t,X_{t}^{i},\mathcal{L}(X_{t}^{i}))\eta_{t}^{i}\mathrm{d}t]=0$, $\phi_{i}(t,X_{t}^{i},\mathcal{L}(X_{t}^{i}))\geq 0$ and $\eta_{t}^{i}\geq 0$ for $i=1,2$. Similarly, we have $\mathbb{E}[\int_{0}^{T}G(t)\bar{Y}_{t}\mathrm{d}\bar{\zeta}_{t}]\leq 0$. Exploiting Assumption 3, it holds that
$0$ $\leq\mathbb{E}\Big[\left(\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{1}),Y_{t}^{2},Z_{t}^{2})-\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2})\right)\left(\bar{X}_{t},\bar{Y}_{t},\bar{Z}_{t}\right)$
$\qquad\quad-\beta\left(|\bar{X}_{t}|^{2}+|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2}\right)\Big].$
However, it follows from (A3) in Assumption 1 that
$\left(\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{1}),Y_{t}^{2},Z_{t}^{2})-\Phi(t,X_{t}^{2},\mathcal{L}(X_{t}^{2}),Y_{t}^{2},Z_{t}^{2})\right)(\bar{X}_{t},\bar{Y}_{t},\bar{Z}_{t})$
$\qquad\leq\frac{M}{2}\mathcal{W}_{2}(\mathcal{L}(X_{t}^{1}),\mathcal{L}(X_{t}^{2}))\left(|X_{t}^{1}-X_{t}^{2}|+|Y_{t}^{1}-Y_{t}^{2}|+|Z_{t}^{1}-Z_{t}^{2}|\right)$
$\qquad\leq\frac{M}{2}\left(|X_{t}^{1}-X_{t}^{2}|^{2}+|Y_{t}^{1}-Y_{t}^{2}|^{2}+|Z_{t}^{1}-Z_{t}^{2}|^{2}\right).$
Hence, combining the above two inequalities leads to $\bar{X}=0,\bar{Y}=0$ and $\bar{Z}=0$. It then follows that $\bar{\alpha}=0$ and $\int_{0}^{\cdot}G(s)\mathrm{d}\bar{\zeta}_{s}=0$. Moreover, we have $\int_{0}^{t}A(s)\bar{\eta}_{s}\mathrm{d}s+\int_{0}^{t}B(s)\mathbb{E}\left[\bar{\eta}_{s}\right]\mathrm{d}s=0$ for all $t\in[0,T]$. Taking expectations on both sides of the above result, together with (C1) in Assumption 3, we conclude that $\mathbb{E}\left[\bar{\eta}_{t}\right]=0$, for $m$-a.e. $t\in[0,T]$, and hence $\int_{0}^{t}A(s)\bar{\eta}_{s}\mathrm{d}s=0$, for $m$-a.e. $t\in[0,T]$. This indicates that $\bar{\eta}=0$ in $\mathbb{L}_{\mathbb{F}}^{2}$ by utilizing Assumption 3-(C1). ∎
Remark 5.5.
Here, we only consider the case where $Y_{T}=g(X_{T},\mu_{T})$ with $g=0$ in the constrained FBSDE (5)-(27). However, the uniqueness result still holds if we additionally assume:
  • (C5)
    Set $\Psi(x,\mu):=\nabla_{x}g(x,\mu)+\int_{\mathbb{R}^{n}}\partial_{\mu}g(x^{\prime},\mu)(x)\mu(\mathrm{d}x^{\prime})$ for $(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$. Then, $\Psi$ satisfies the monotone condition $(\Psi(x_{1},\mu)-\Psi(x_{2},\mu))(x_{1}-x_{2})\geq 0$ and $\int_{\mathbb{R}^{n}}(\Psi(x,\mu_{1})-\Psi(x,\mu_{2}))x^{\prime}(\mu_{1}-\mu_{2})(\mathrm{d}x^{\prime})\geq 0$ for $x,x_{1},x_{2}\in\mathbb{R}^{n}$ and $\mu,\mu_{1},\mu_{2}\in\mathcal{P}_{2}(\mathbb{R}^{n})$.
Finally, we examine the stability of the solution to the constrained FBSDE (5)-(27). To this end, we introduce the following enhanced conditions:
  • (C1’)
    The constrained function $\phi$ is independent of the control variable and admits the linear form that $\phi(t,x,\mu)=A(t)x+B(t)\int_{\mathbb{R}^{n}}x\mu(\mathrm{d}x)+C(t)$ for $(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$. Here, the coefficients $A,B,C\in C([0,T],\mathbb{R}^{n})$ satisfying $|A(t)|\geq c_{0}>0$ and $|A(t)+B(t)|\geq c_{0}>0$ for some $c_{0}>0$ independent of $t\in[0,T]$.
Then, we have
Theorem 5.6.
Let Assumption 2, (C1’), (C3)-(C4) of Assumption 3 hold and $\xi_{1},\xi_{2}\in L^{2}((\Omega,\mathcal{F}_{0},\mathbb{P});\mathbb{R}^{n})$. Denote by $(\widehat{X}^{i},\widehat{\alpha}_{t}^{i},\widehat{\zeta}^{i}_{t},Y^{i},Z^{i},\eta^{i})$ the unique solutions to the FBSDE (5)-(27) with respective initial data $\xi_{i}$. Set $\bar{X}_{t}:=X_{t}^{1}-X_{t}^{2}$, $\bar{\zeta}_{t}:=\zeta_{t}^{1}-\zeta_{t}^{2}$, $\bar{Y}_{t}:=Y_{t}^{1}-Y_{t}^{2}$, $\bar{Z}_{t}:=Z_{t}^{1}-Z_{t}^{2}$ and $\bar{\eta}_{t}:=\eta_{t}^{1}-\eta_{t}^{2}$ for $t\in[0,T]$. Then, there exists a constant $C>0$ depending on $(T,M,\alpha,\beta)$ only such that
$\mathbb{E}\left[\left|\bar{X}_{T}\right|^{2}+\int_{0}^{T}\left(|\bar{X}_{t}|^{2}+|\bar{\alpha}_{t}|^{2}+|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2}+|\bar{\eta}_{t}|^{2}\right)\mathrm{d}t+\left|\int_{0}^{T}G(t)\mathrm{d}\bar{\zeta}_{t}\right|^{2}\right]$
$\qquad\qquad\leq\mathbb{E}\left[|\xi_{1}-\xi_{2}|^{2}\right].$ (31)
Proof.
Recall $\Delta b(t),\Delta\sigma(t),\Delta F(t)$ defined in the proof of Theorem 5.4. Then, one has
$\int_{0}^{t}G(s)\mathrm{d}\bar{\zeta}_{s}=\bar{X}_{t}-(\xi_{1}-\xi_{2})-\int_{0}^{t}\Delta b(s)\mathrm{d}s-\int_{0}^{t}\Delta\sigma(s)\mathrm{d}s.$ (32)
On the other hand, applying Itô’s formula to $|\bar{X}_{t}|^{2}$ leads to
$\mathrm{d}|\bar{X}_{t}|^{2}=2\left(\Delta b(t)^{\top}\bar{X}_{t}+\mathrm{tr}\left(\Delta\sigma\Delta\sigma^{\top}(t)\right)\right)\mathrm{d}t+2\bar{X}_{t}^{\top}\Delta\sigma(t)\mathrm{d}W_{t}+2\bar{X}_{t}^{\top}G(t)\mathrm{d}\bar{\zeta}_{t}.$ (33)
By the Young’s inequality, $\left|\int_{0}^{t}2\bar{X}_{s}^{\top}G(s)\mathrm{d}\bar{\zeta}_{s}\right|\leq\epsilon\sup_{0\leq s\leq t}|\bar{X}_{s}|^{2}+\frac{1}{\epsilon}\left|\int_{0}^{t}G(s)\mathrm{d}\bar{\zeta}_{s}\right|^{2}$ for any $\epsilon\in(0,1)$. Plugging the above inequality into the equality (33), and taking expectations on both sides, we arrive at, for any $t\in[0,T]$,
$\mathbb{E}\left[\sup_{0\leq s\leq t}|\bar{X}_{s}|^{2}\right]\leq\mathbb{E}\left[\sup_{0\leq s\leq t}\int_{0}^{s}2\left(\Delta b(r)^{\top}\bar{X}_{r}+\mathrm{tr}\left(\Delta\sigma\Delta\sigma^{\top}(r)\right)\right)\mathrm{d}r\right]$
$\qquad+\mathbb{E}\left[\sup_{0\leq s\leq t}\int_{0}^{s}2\bar{X}_{r}^{\top}\Delta\sigma(r)\mathrm{d}W_{r}\right]+\epsilon\mathbb{E}\left[\sup_{0\leq s\leq t}|\bar{X}_{s}|^{2}\right]+\frac{1}{\epsilon}\mathbb{E}\left[\left|\int_{0}^{t}G(s)\mathrm{d}\bar{\zeta}_{s}\right|^{2}\right].$
Using (C4) of Assumption 3, the equality (32), the BDG’s inequality and the Gronwall’s inequality, we deduce the existence of a constant $C>0$ depending on $(M,\epsilon,T)$ only such that
$\mathbb{E}\left[\sup_{0\leq t\leq T}|\bar{X}_{t}|^{2}\right]\leq C\mathbb{E}\left[|\xi^{1}-\xi^{2}|^{2}+\int_{0}^{T}(|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2})\mathrm{d}t\right].$ (34)
With the help of the BSDE in (5), we have, for all $t\in[0,T]$,
$\int_{t}^{T}A(s)\bar{\eta}_{s}\mathrm{d}s+\int_{t}^{T}B(s)\mathbb{E}\left[\bar{\eta}_{s}\right]\mathrm{d}s=\bar{Y}_{T}-Y_{t}+\int_{t}^{T}\Delta F(s)\mathrm{d}s-\int_{t}^{T}\bar{Z}_{s}\mathrm{d}W_{s}.$ (35)
Applying the Itô’s formula to $|\bar{Y}_{t}|^{2}$, we deduce, for all $t\in[0,T]$,
$\left|\bar{Y}_{t}\right|^{2}+\int_{t}^{T}|\bar{Z}_{s}|^{2}\mathrm{d}s$ $=2\int_{t}^{T}\bar{Y}_{s}^{\top}\Delta F(s)\mathrm{d}s-2\int_{t}^{T}\bar{Y}_{s}\bar{Z}_{s}\mathrm{d}W_{s}-2\int_{t}^{T}\bar{\eta}_{s}A(s)^{\top}\bar{Y}_{s}\mathrm{d}s$
$\quad-2\int_{t}^{T}\mathbb{E}\left[\bar{\eta}_{s}\right]B(s)^{\top}\bar{Y}_{s}\mathrm{d}s.$
By using the Young’s inequality again, for all $t\in[0,T]$,
$\left|2\int_{t}^{T}\bar{\eta}_{s}A(s)^{\top}\bar{Y}_{s}\mathrm{d}s\right|$ $\leq\epsilon\sup_{t\leq s\leq T}|\bar{Y}_{s}|^{2}+\frac{1}{\epsilon}\left|\int_{t}^{T}A(s)^{\top}\bar{\eta}_{s}\mathrm{d}s\right|^{2},$
$\left|2\int_{t}^{T}\mathbb{E}\left[\bar{\eta}_{s}\right]B(s)^{\top}\bar{Y}_{s}\mathrm{d}s\right|$ $\leq\epsilon\sup_{t\leq s\leq T}|\bar{Y}_{s}|^{2}+\frac{1}{\epsilon}\left|\int_{t}^{T}B(s)^{\top}\mathbb{E}\left[\bar{\eta}_{s}\right]\mathrm{d}s\right|^{2}.$
Using Assumption 3-(C4), the condition (C1’), the BDG’s inequality and the Gronwall’s inequality, we can similarly derive, for some constant $C>0$ depending on $(M,\epsilon,T)$ only,
$\mathbb{E}\left[\sup_{0\leq t\leq T}|\bar{Y}_{t}|^{2}+\int_{0}^{T}|\bar{Z}_{t}|^{2}\mathrm{d}t\right]\leq C\mathbb{E}\left[|\bar{X}_{T}|^{2}+\int_{0}^{T}|\bar{X}_{t}|^{2}\mathrm{d}t\right].$ (36)
Applying integration by parts to $\bar{X}_{t}^{\top}\bar{Y}_{t}$ from $t=0$ to $T$, we have
$\bar{X}_{T}\bar{Y}_{T}$ $=\int_{0}^{T}(-\Delta\bar{F}(t)\bar{X}_{t}+\bar{Y}\Delta\bar{b}(t)+\mathrm{tr}(\bar{Z}_{t}\Delta\bar{\sigma}(t)))\mathrm{d}t+\int_{0}^{T}(\bar{X}_{t}\bar{Z}_{t}+\bar{Y}_{t}\Delta\sigma)\mathrm{d}W_{t}$
$\quad+\int_{0}^{T}\left\{A(t)\bar{\eta}_{t}+B(t)\mathbb{E}[\bar{\eta}_{t}]\right\}\bar{X}_{t}+\int_{0}^{T}G(t)\bar{Y}_{t}\mathrm{d}\bar{\zeta}_{t}.$
Taking expectations on both sides of the above equality and following the proof of Theorem 5.4, we deduce that
$-\mathbb{E}\left[\bar{Y}_{0}^{\top}\left(\xi^{1}-\xi^{2}\right)\right]+\beta\mathbb{E}\left[\int_{0}^{T}(|\bar{X}_{t}|^{2}+|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2})\mathrm{d}t\right]\leq 0.$ (37)
As a result, the inequality (37) leads to that
$\alpha\mathbb{E}\left[\left|\bar{X}_{T}\right|^{2}\right]+\beta\mathbb{E}\left[\int_{0}^{T}(|\bar{X}_{t}|^{2}+|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2})\mathrm{d}t\right]$ $\leq\mathbb{E}\left[\bar{Y}_{0}^{\top}(\xi^{1}-\xi^{2})\right]\leq\frac{2}{\epsilon}\mathbb{E}\left[\left|\bar{Y}_{0}\right|^{2}\right]$
$\quad+\frac{1}{2\epsilon}\mathbb{E}\left[\left|\xi^{1}-\xi^{2}\right|^{2}\right].$
Choosing $\epsilon>0$ small enough and taking into account of (34) and (36), we can find a constant $C>0$ depending on $M,T,\alpha,\beta$ only such that
$\mathbb{E}\left[\left|\bar{X}_{T}\right|^{2}+\int_{0}^{T}\left(|\bar{X}_{t}|^{2}+|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2}\right)\mathrm{d}t\right]\leq C\mathbb{E}\left[\left|\xi^{1}-\xi^{2}\right|^{2}\right].$ (38)
Combining the Lipschitz continuity of $\hat{u}:[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{n}\times\mathbb{R}^{n\times r}\mapsto U\subset\mathbb{R}^{l}$ in Assumption 3-(C3), (32), (35) and the condition (C1’), we conclude the stability (5.6) from (38) above. ∎
Remark 5.7.
When considering $Y_{T}=g(X_{T},\mu_{T})$, we further assume that
  • (C5’)
    Set $\Psi(x,\mu):=\nabla_{x}g(x,\mu)+\int_{\mathbb{R}^{n}}\partial_{\mu}g(y,\mu)(x)\mu(\mathrm{d}y)$ for $(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$. The function $\Psi$ is Lipschitz continuous and satisfies the coercivity condition that, for some $\alpha,M>0$ and $(x^{1},\mu^{1},X^{1}),(x^{2},\mu^{2},X^{2})\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times L^{2}((\Omega,\mathcal{F}_{T},\mathbb{P}),\mathbb{R}^{n})$,
    $|\Psi(x^{1},\mu^{1})-\Psi(x^{2},\mu^{2})|\leq M(|x^{1}-x^{2}|+\mathcal{W}_{2,\mathbb{R}^{n}}(\mu^{1},\mu^{2})),$
    $\mathbb{E}\left[\left\langle\Psi(X^{1},\mathcal{L}(X^{1}))-\Psi(X^{2},\mathcal{L}(X^{2})),X^{1}-X^{2}\right\rangle\right]\geq\alpha\mathbb{E}[|X^{1}-X^{2}|^{2}].$
In addition to the setting in Theorem 5.6, let $Y^{i}$ have terminal value $g^{i}(X_{T}^{i},\mu_{T}^{i})$ and $g^{1},g^{2}\in C^{1;1}(\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n});\mathbb{R})$ with $g^{1}$ satisfying the condition (C5’). Then, it holds that
$\mathbb{E}\left[\left|\bar{X}_{T}\right|^{2}+\int_{0}^{T}\left(|\bar{X}_{t}|^{2}+|\bar{\alpha}_{t}|^{2}+|\bar{Y}_{t}|^{2}+|\bar{Z}_{t}|^{2}+|\bar{\eta}_{t}|^{2}\right)\mathrm{d}t+\left|\int_{0}^{T}G(t)\mathrm{d}\bar{\zeta}_{t}\right|^{2}\right]$
$\qquad\qquad\qquad\leq C\left\{\mathbb{E}\left[|\xi_{1}-\xi_{2}|^{2}\right]+\sup_{(x,\mu)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})}\left|\Psi^{1}(x,\mu)-\Psi^{2}(x,\mu)\right|^{2}\right\}$
with $\Psi^{i}(x,\mu):=\nabla_{x}g^{i}(x,\mu)+\int_{\mathbb{R}^{n}}\partial_{\mu}g^{i}(y,\mu)(x)\mu(\mathrm{d}y)$ for $i=1,2$.
Acknowledgements L. Bo and J. Wang are supported by National Natural Science of Foundation of China (No. 12471451), Natural Science Basic Research Program of Shaanxi (No. 2023-JC-JQ-05), Shaanxi Fundamental Science Research Project for Mathematics and Physics (No. 23JSZ010). X. Yu and J. Wang are supported by the Hong Kong RGC General Research Fund (GRF) under grant No. 15214125 and by the Hong Kong Polytechnic University research grant under No. P0045654.
Appendix: Proofs of Auxiliary Results
This appendix collects proofs of the auxiliary results presented in the main body of the paper.
Proof of Lemma 3.4.
For any $\omega\in\Omega$, define the set $K(t,\omega):=K(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}^{\alpha,\zeta}))$ with $(t,\omega)\in[0,T]\times\Omega$. Since $K(t,\omega)$ is convex and closed, it follows that
$\bigg($ $\int_{U}b(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),u)q_{t}(\omega,\mathrm{d}u),\int_{U}\sigma\sigma^{\top}(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),u)q_{t}(\omega,\mathrm{d}u),$
$\quad\int_{U}f(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),u)q_{t}(\omega,\mathrm{d}u),\int_{U}\phi^{i}(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),u)q_{t}(\omega,\mathrm{d}u),~i\in I\bigg)\in K(t,\omega)$
is equivalent to
$U(t,\omega):=\Big\{u\in U;$~ $\int_{U}b(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),u)q_{t}(\omega,\mathrm{d}u)=b(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),u),$
$\int_{U}\sigma\sigma^{\top}(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),u)q_{t}(\omega,\mathrm{d}u)=\sigma\sigma^{\top}(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),u),$
$\int_{U}f(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),u)q_{t}(\omega,\mathrm{d}u)\geq f(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),u),$
$\int_{U}\phi^{i}(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),u)q_{t}(\omega,\mathrm{d}u)\leq\phi^{i}(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),u),~i\in I\Big\}\neq\varnothing.$
[…]
Next, we show that $\alpha=(\alpha_{t})_{t\in[0,T]}$ is an admissible (strict) control. Using the growth condition imposed on $f$ in Assumption 1-(A6), we have from the above construction of $\alpha$ that
$+\infty$ $>\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}f(t,X_{t},\mathcal{L}^{R}(X_{t}),u)q_{t}(\mathrm{d}u)\mathrm{d}t\right]\geq\mathbb{E}^{R}\left[\int_{0}^{T}f(t,X_{t},\mathcal{L}^{R}(X_{t}),\alpha_{t})\mathrm{d}t\right]$
$\geq-\int_{0}^{T}M(1+\mathbb{E}^{R}[|X_{t}|^{p}]+M_{2}(R^{X}(t)))\mathrm{d}t+M_{1}\mathbb{E}^{R}\left[\int_{0}^{T}|\alpha_{t}|^{p}\mathrm{d}t\right].$
This, together with Lemma 3.2, indicates that $\mathbb{E}^{R}[\int_{0}^{T}|\alpha_{t}|^{p}\mathrm{d}t]<\infty$. On the other hand, using the construction of $\alpha$ again, we deduce that, for any $g\in C_{b}^{2}(\mathbb{R}^{n})$ and $(t,\omega)\in[0,T]\times\Omega$,
$g(X_{t}(\omega))-\int_{0}^{t}\mathbb{G}g(s,X_{s}(\omega),\mathcal{L}^{R}(X_{s}),\alpha_{s}(\omega))\mathrm{d}s-\int_{0}^{t}\nabla_{x}g(X_{s}(\omega))G(s)\mathrm{d}\zeta_{s}(\omega)$
$\qquad-\sum_{0<s\leq t}\left(g(X_{s}(\omega))-g(X_{s-}(\omega))-\nabla_{x}g(X_{s-}(\omega))\Delta X_{s}(\omega)\right)=M_{t}^{R}g(\omega)$
is an $(R,\mathbb{F})$-martingale as $R$ is a control rule in Definition 3.1. Moreover, we have, for $m\otimes R$-$\mathrm{a.s.}~(t,\omega)\in[0,T]\times\Omega$,
$\phi^{i}\left(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),\alpha_{t}(\omega)\right)\geq\int_{U}\phi^{i}(t,X_{t}(\omega),\mathcal{L}^{R}(X_{t}),u)q_{t}(\omega,\mathrm{d}u)\geq 0,~~\forall i\in I.$
This yields that $\alpha=(\alpha_{t})_{t\in[0,T]}$ is admissible. Furthermore, by the construction of $\alpha$, it holds that
$J(\alpha,\zeta)$ $=\mathbb{E}^{R}\left[\int_{0}^{T}f(t,X_{t},\mathcal{L}^{R}(X_{t}),\alpha_{t})\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}\right]$
$\leq\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}f(t,X_{t},R^{X}(t),u)q_{t}(\mathrm{d}u)\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}\right]=\Gamma(R).$
Thus, we complete the proof of the lemma. ∎
Proof of Lemma 3.7.
[…]
$\tilde{A}_{i}^{R}:=\left\{\omega\in\Omega;~\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t\geq 0,~\forall 0\leq t_{1}<t_{2}\leq T\right\}.$
We first show that $m\otimes R(A_{i}^{R})=T$ iff $R(\tilde{A}_{i}^{R})=1$. Here, recall that $A_{i}^{R}\subset[0,T]\times\Omega$ is defined by (5).
Necessity. For any $\omega\in\Omega$, define the $\omega$-section of $A_{i}^{R}$ that
$A_{i}^{R}(\omega)=\left\{t\in[0,T];~\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\geq 0\right\}.$
Then, it follows from Fubini’s Theorem that $T=m\otimes R(A_{i}^{R})=\int_{\Omega}m(A_{i}^{R}(\omega))R(\mathrm{d}\omega)$. This implies that there exists some $R$-null set $N\in\mathcal{F}$ such that $m(A_{i}^{R}(\omega))=T$ for all $\N\omega\in\Omega\backslash$ N. For such $\omega$ and $0\leq t_{1}<t_{2}\leq T$, it holds that
$\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t$
$\qquad=\int_{(t_{1},t_{2}]\cap A_{i}^{R}(\omega)}\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t\geq 0.$
Consequently, $\omega\in\tilde{A}_{i}^{R}$, or more precisely, $\Omega\backslash N\subset\tilde{A}_{i}^{R}$, and hence $R(\tilde{A}_{i}^{R})=1$.
Sufficiency. For $\omega\in\tilde{A}_{i}^{R}$, we have $m(A_{i}^{R}(\omega))=T$, and Fubini’s Theorem results in $m\otimes R(A_{i}^{R})=\int_{\Omega}m(A_{i}^{R}(\omega))R(\mathrm{d}\omega)=T$.
[…]
$\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,\psi_{t},R^{X}(t),u)q_{t}(\mathrm{d}u)\mathrm{d}t\geq\limsup_{\ell\to\infty}\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,\psi^{\ell}_{t},R^{X}(t),u)q^{\ell}_{t}(\mathrm{d}u)\mathrm{d}t\geq 0,$
where we utilized the fact that $\phi^{i}$ is u.s.c. in $(x,\mu,u)\in\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$ imposed in Assumption 1-(A2). Hence, $\omega\in\tilde{A}_{i}^{R}$, which yields the closedness of $\tilde{A}_{i}^{R}$.
[…]
$\qquad\int_{t_{1}}^{t_{2}}\left.\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t<0,~\exists 0\leq t_{1}<t_{2}\leq T\right\}:=B_{\ell}.$
Note that $\phi^{i}$ is uniformly u.s.c. in $\mu$ w.r.t. $(t,x,u)\in[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}$, which implies that
$\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)\geq\limsup_{\ell\to\infty}\phi^{i}(t,X_{t}(\omega),R_{\ell}^{X}(t),u)$
uniformly w.r.t. $(t,\omega,u)\in[0,T]\times\Omega\times\mathbb{R}^{l}$, and hence
$\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t\geq\limsup_{\ell\to\infty}\int_{t_{1}}^{t_{2}}\int_{U}\phi^{i}(t,X_{t}(\omega),R_{\ell}^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t$
uniformly w.r.t. $\omega\in\Omega$. Hence, $B_{\ell}$ is an empty set whenever $\ell$ is sufficiently large. Thus, it holds that
$\limsup_{\ell\to\infty}R_{\ell}(\tilde{A}_{i}^{R_{\ell}})$ $\leq\limsup_{\ell\to\infty}\left(R_{\ell}(\tilde{A}_{i}^{R_{\ell}})-R_{\ell}(\tilde{A}_{i}^{R})\right)+\limsup_{\ell\to\infty}R_{\ell}(\tilde{A}_{i}^{R})$
$=\limsup_{\ell\to\infty}R_{\ell}(\tilde{A}_{i}^{R_{\ell}}\backslash\tilde{A}_{i}^{R})+\limsup_{\ell\to\infty}R_{\ell}(\tilde{A}_{i}^{R})\leq R(\tilde{A}_{i}^{R}).$
For the last inequality, we used the fact that $\tilde{A}_{i}^{R}$ is closed and $R_{\ell}$ converges weakly to $R$ as $\ell\to\infty$. This yields that $R(\tilde{A}_{i}^{R})=1$, and hence $m\otimes R(\tilde{A}_{i}^{R})=T$, which completes the proof. ∎